### v5

# Microscopic theory of turbulent mixing: discrete shell structures in scalar concentration

Abstract
We present an exact analytic solution for the one-point distribution of a passive scalar in decaying homogeneous turbulence, in the limit of vanishing viscosity and diffusivity at fixed Schmidt number. The velocity statistics are governed by the Euler ensemble, obtained previously as a spontaneously stochastic solution of the loop equation derived from the Navier–Stokes equations in the extreme turbulent limit. Here we extend the formulation to the liquid (Lagrangian) loop, advected with the flow, and show that the Euler ensemble solves this equation as well. The advection–diffusion problem for the scalar is recast exactly as a closed linear loop equation and solved by loop calculus.
For a localized initial condition, the solution consists of a sequence of expanding concentric shells: the radial scalar profile is quantized and piecewise parabolic, with gaps organized by Euler totients—an arithmetic structure not anticipated by conventional models. This shell pattern is the unique solution within the Euler ensemble; finite diffusivity or sustained forcing smooths it. The result provides a geometric skeleton for scalar transport in extreme decaying turbulence and may be relevant in astrophysical or quantum-fluid regimes where dissipation is negligible. These findings offer a parameter-free benchmark for scalar mixing in the high–Reynolds-number limit.
keywords:
Turbulence, Fractal, Transport Equation, Passive Scalar, Fixed Point, Velocity Circulation, Loop Equations
1 Preface: Microscopic theory of decaying turbulence
Turbulence remains one of the oldest unsolved problems in classical physics. As Feynman noted in his 1964 Lectures, the analysis of turbulent flows has resisted a complete mathematical formulation for over a century. Despite extensive theoretical work and direct numerical simulations (DNS), a predictive, analytically tractable solution has been elusive.
This paper presents an exact solution for passive-scalar mixing in decaying turbulence. It builds on a recent solution of the velocity problem: an exact description of decaying turbulence in the vanishing-viscosity limit, derived from first principles via loop calculus. In that formulation, the Navier–Stokes dynamics are recast in loop space, and a dual description emerges in which the turbulent regime is captured by a spontaneously stochastic probability measure supported on exact Navier–Stokes solutions — the Euler ensemble. Discrete structures arise in momentum-loop space in the extreme turbulent limit, with a regulator $N\to\infty$ and fixed “turbulent viscosity” $\tilde{\nu}=\nu N^{2}$.
Three conceptual shifts underlie the approach:
  1. 1.
    Loop-space reformulation. Rather than tracking the velocity field as a function of $(\mathbf{r},t)$, we study the evolution of circulation along closed loops in physical space. This yields a closed, linear equation for the loop functional—the analogue of a Schrödinger equation in loop space.
  2. 2.
    Spontaneous stochasticity at $\nu\to 0$. The turbulent regime is described by a stochastic solution that selects a discrete ensemble of momentum loops (realized as random walks on regular star polygons). The solution is exact in the limit $\nu\to 0$ with $\tilde{\nu}$ fixed, where a smooth loop is replaced by an $N$-gon and $N\to\infty$.
  3. 3.
    Static and liquid loop equations. The classical (Eulerian) loop equation for a fixed loop has diffusion and advection terms. We also formulate the Lagrangian (“liquid”) loop equation for a loop advected by the flow; by Kelvin’s theorem the advection term cancels and only the diffusion operator remains. We show that both equations are solved by the same turbulent attractor—the Euler ensemble.
  4. 4.
    Arithmetic structure. The resulting statistics exhibit an arithmetic organization (Euler totients and zeta sums) that will surface in the passive-scalar solution below.
DNS performed by Sreenivasan and collaborators already supports several predictions of the Euler ensemble solution in decaying turbulence. The predicted decay of turbulent kinetic energy, $\mathrm{TKE}(t)\sim t^{-5/4}$, agrees well with experiment and DNS, unlike the classical Kolmogorov–Saffman model; further agreement is seen in refined statistics such as effective index obtained by log derivative of the second moment of velocity $\left\langle(\boldsymbol{v}(\boldsymbol{r})-\boldsymbol{v}(0))^{2}\right\rangle_{v}$ as a function of the scaling variable $\log|\boldsymbol{r}|/\sqrt{t}$
Here we extend the velocity solution to a second fundamental question: the transport of a passive scalar in decaying turbulence. Scalar mixing (temperature, concentration) is central in environmental, industrial and astrophysical flows. Conventional treatments predict Gaussian spreading of a localized scalar, smoothed by molecular diffusion and eddy transport, but rely on surrogate velocity statistics or closure assumptions and lose predictive power at high Reynolds number.
The passive scalar does not feed back on the flow; it is advected by the velocity field and diffused molecularly. Mathematically, one seeks the scalar density as a functional of the velocity, then averaged over its statistics. In the classical literature this was solved exactly only in surrogate models (e.g. the Gaussian, $\delta$-correlated Kraichnan model Falkovich et al. (2001)). By contrast, the loop formulation yields a closed linear equation for the scalar loop functional that we solve exactly at fixed Schmidt number in the decaying regime.
We find that, in the Euler ensemble, a localized initial condition produces a hierarchy of expanding concentric shells: the radial scalar profile is quantized and piecewise parabolic, with gaps organized by Euler totients. This “shell” pattern is the geometric skeleton of scalar transport in ideal decaying turbulence; finite diffusivity or sustained forcing smooths it. The result is a parameter-free, analytic solution for passive-scalar statistics in the extreme turbulent limit.
Paper outline
Section 2 states the governing equations and the decaying-regime setting. Section 3 reviews loop calculus and the static/liquid loop equations, and shows their equivalence on the Euler ensemble in the extreme turbulent limit. Section 4 derives the passive-scalar loop equation and presents its exact solution, including the shell structure. Section 5 discusses the physical picture in loop space and compares the Euler-ensemble solution with established phenomenologies (Kolmogorov, Heisenberg, multifractal). Conclusions and implications (including number-theoretic structure, possible relevance in quantum fluids or astrophysical flows, and directions for DNS validation) are summarized in Section 6.
2 Introduction to the turbulent mixing
Turbulent mixing—the transport and diffusion of a passive scalar by a turbulent velocity field—is a fundamental process in many natural and technological systems. It governs the dispersion of pollutants, the distribution of heat and salinity in the oceans, and mixing in stars and industrial flows Sreenivasan (2019b); Dimotakis (2005); Warhaft (2000). The governing equations are the incompressible Navier–Stokes equations coupled to the scalar advection–diffusion equation:
$\frac{\partial\boldsymbol{v}}{\partial t}+(\boldsymbol{v}\cdot\nabla)\boldsymbol{v}=-\nabla p+\nu\nabla^{2}\boldsymbol{v},$ (1a)
$\nabla\cdot\boldsymbol{v}=0,$ (1b)
$\frac{\partial T}{\partial t}+(\boldsymbol{v}\cdot\nabla)T=D\nabla^{2}T,$ (1c)
where $\boldsymbol{v}(\boldsymbol{r},t)$ is the velocity field, $p$ is the pressure, $T(\boldsymbol{r},t)$ is the passive scalar, $\nu$ is the kinematic viscosity, and $D$ is the scalar diffusivity. The dimensionless Schmidt number is $\mathrm{Sc}=\nu/D$. Many theoretical, experimental, and numerical studies have addressed passive-scalar behavior, most often through global observables such as variance decay, spectra, or correlation functions (see, e.g., the review Falkovich et al. (2001)). A key open question concerns the spatial structure that emerges from a localized initial condition in decaying turbulence. Traditional treatments suggest rapid Gaussian spreading under turbulent advection and molecular diffusion, but a closed-form solution for this regime has not been available.
We present an exact solution for the one-point distribution of a passive scalar advected by a velocity field drawn from the Euler ensemble—a spontaneously stochastic probability measure supported on solutions of the unforced Navier–Stokes equations in the vanishing-viscosity limit Migdal (2024). Using loop calculus, we solve the linear advection–diffusion equation without surrogate statistics or closure assumptions. The resulting scalar field exhibits a sequence of concentric spherical shells with quantized, piecewise-parabolic radial profiles and a smooth central value. This sharply ordered structure is a deterministic prediction of the ensemble and persists in the extreme turbulent limit; finite diffusivity smooths it. The effect may be relevant in regimes where dissipation is negligible, such as certain astrophysical or quantum-fluid settings.
3 Decaying turbulence as stochastic motion of liquid loops
The recent solution of decaying turbulence Migdal (2023, 2024, 2025a) follows a microscopic route: we start from the unforced Navier-Stokes equations and work directly with their associated loop dynamics. We make no modelling substitutions (no Gaussian surrogates, no closure ansätze); the conclusions recorded here follow from the unforced dynamics in the decaying regime rather than from auxiliary hypotheses. The nontrivial structures we report are thus features of the equations, not of the modelling. In this section we present the ideas step by step, highlighting both the mathematical derivation and the physical interpretation of the resulting picture.
3.1 The Euler and the Lagrange pictures in the circulation dynamics
There are two complementary descriptions of the dynamics: the Eulerian picture, which studies the distribution of the velocity field in a rest frame, and the Lagrangian picture, which follows material points of the flow. We use both within loop dynamics.
[…]
$\Gamma_{C}[v]=\int_{C}d\mathbf{r}\cdot\mathbf{v}(\mathbf{r})=\oint d\theta\ \mathbf{C}^{\prime}(\theta)\cdot\mathbf{v}(\mathbf{C}(\theta)).$ (2)
Here $C$ is fixed in space while the flow passes through it; the circulation measures the vorticity flux through the surface bounded by $C$. Its time derivative is
$\partial_{t}\Gamma_{C}[v]=\oint d\theta\ \mathbf{C}^{\prime}(\theta)\cdot\big{(}-\nu\,\nabla\times\boldsymbol{\omega}(\mathbf{C}(\theta))\;-\;\boldsymbol{\omega}(\mathbf{C}(\theta))\times\mathbf{v}(\mathbf{C}(\theta))\big{)},$ (3)
where the remaining Navier-Stokes terms reduce to a gradient and vanish upon integration over the closed loop:
$\oint d\theta\ \mathbf{C}^{\prime}(\theta)\cdot\nabla\!\left(p+\tfrac{1}{2}\,\mathbf{v}^{2}\right)=0.$
In the Lagrangian picture each point of the loop moves with the local velocity, which adds a transport term by the chain rule:
$\frac{D}{Dt}\Gamma_{C}[v]=\partial_{t}\Gamma_{C}[v]+\oint d\theta\ \frac{\delta\Gamma_{C}[v]}{\delta\mathbf{C}(\theta)}\cdot\mathbf{v}(\mathbf{C}(\theta)).$ (4)
Using Stokes’ theorem one finds
$\frac{\delta\Gamma_{C}[v]}{\delta\mathbf{C}(\theta)}=\mathbf{C}^{\prime}(\theta)\times\boldsymbol{\omega}(\mathbf{C}(\theta)),$ (5)
so that the convective term cancels the corresponding term above:
$\frac{D}{Dt}\Gamma_{C}[v]=\oint d\theta\ \mathbf{C}^{\prime}(\theta)\cdot\big{(}-\nu\,\nabla\times\boldsymbol{\omega}(\mathbf{C}(\theta))\;-\;\cancel{\boldsymbol{\omega}(\mathbf{C}(\theta))\times\mathbf{v}(\mathbf{C}(\theta))}\big{)}.$ (6)
This Kelvin cancellation implies that only the viscous contribution remains; in the formal limit $\nu\to 0$ the circulation is conserved.
In the physical world, the limit $\nu\to 0$ in the Navier-Stokes flow is not the Euler flow but the extreme turbulent regime, in which strong vorticity gradients compensate the small viscosity to produce a nontrivial limit—the setting of decaying turbulence that we study.
The static and liquid loop equations thus describe the same physics, and—as shown below—are both solved by the same turbulent attractor, the Euler ensemble Migdal (2023). Precisely, the Euler ensemble solves the liquid loop equation, and the advection term $\mathbf{v}\times\boldsymbol{\omega}$ in the static equation vanishes on this solution.
3.2 The new loop calculus
We now recall the loop calculus Makeenko & Migdal (1979); Migdal (1983, 1998), reviewed with improvements in Migdal (2025b). The basic object is the area derivative, obtained from Stokes’ theorem for the first variation of the circulation:
$\delta\Gamma_{C}[v]=\oint d\theta\ \mathbf{C}^{\prime}(\theta)\times\delta\mathbf{C}(\theta)\cdot\boldsymbol{\omega}(\mathbf{C}(\theta)).$ (7)
For a parametrization-invariant functional $\Phi[C]$, the area derivative $\delta\Phi[C]/\delta\boldsymbol{\sigma}(\theta)$ is defined by
$\delta\Phi[C]=\oint d\theta\ \mathbf{C}^{\prime}(\theta)\times\frac{\delta\Phi[C]}{\delta\boldsymbol{\sigma}(\theta)}\cdot\delta\mathbf{C}(\theta).$ (8)
Functionals with a first variation of this form are called Stokes-type functionals Makeenko & Migdal (1979); Migdal (1983); any function of circulation is such a functional, as is the Wilson loop in gauge theory and the area of a minimal surface spanned by $C$ (but not the perimeter).
A convenient expression relates the area derivative to the left/right functional derivatives with respect to $\mathbf{C}^{\prime}(\theta\pm 0)$:
$\frac{\delta\Phi[C]}{\delta\boldsymbol{\sigma}(\theta)}=\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta-0)}\times\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta+0)}\,\Phi[C].$ (9)
Differentiating (8) and using integration by parts with periodicity gives
$\frac{\delta}{\delta\mathbf{C}(\theta^{\prime})}=-\partial_{\theta^{\prime}}\,\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta^{\prime})},$ (10)
and integrating the delta function across $(\theta-0,\theta+0)$ yields the stated relation (only the $\theta\pm 0$ terms survive in the cross product). Similarly, the local gradient at a point on the loop is the difference of the left/right derivatives:
$\frac{\partial}{\partial\mathbf{C}(\theta,t)}=\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta-0)}-\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta+0)}.$ (11)
The functional derivative $\delta\Phi[C]/\delta\mathbf{C}^{\prime}(\theta)$ is defined modulo an additive constant vector, but this ambiguity does not affect (11). It is also invariant under reparametrizations $\theta\mapsto f(\theta)$, $f^{\prime}(\theta)>0$. Moreover, any Stokes-type functional obeys
$\frac{\partial}{\partial\mathbf{C}(\theta)}\Phi[C]=0\quad\forall\,\theta,$ (12)
since the integration region in (8) is infinitesimal while the integrand is finite. Likewise, the area derivative annihilates any local function of $\mathbf{C}(\theta)$:
$\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta-0)}\times\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta+0)}\,F(\mathbf{C}(\theta_{1}))=0,$ (13)
because the cross of two gradients vanishes.
3.3 Observables in the loop dynamics
Given a loop functional, the loop calculus provides a route to local observables, such as velocity correlations or the decay law of turbulent kinetic energy. The static loop functional at fixed $C$ is the characteristic function of the circulation distribution $W_{C}(\Gamma)$:
$\Psi[C,t]=\int_{-\infty}^{\infty}d\Gamma\ W_{C}(\Gamma)\,\exp\left(\frac{\mathrm{i}\,}{\nu}\,\Gamma\right),$ (14)
$W_{C}(\Gamma)=\Big{\langle}\delta\big{(}\Gamma-\Gamma_{C}[v]\big{)}\Big{\rangle}_{v}.$ (15)
This identity is useful in DNS, where velocity statistics are estimated from many realizations in space and time; by isotropy and homogeneity one can further average over translations and rotations of $C$.
The inverse direction—recovering velocity statistics from $\Psi[C,t]$—follows by functional differentiation. For example, a second area derivative at separated points gives the two-point vorticity correlation:
$\Big{\langle}\omega_{i}(\mathbf{C}(\theta_{1}))\,\omega_{j}(\mathbf{C}(\theta_{2}))\,\exp\left(\frac{\mathrm{i}\,}{\nu}\,\Gamma_{C}[v]\right)\Big{\rangle}_{v}=\left(\frac{\nu}{\mathrm{i}\,}\right)^{2}\frac{\delta^{2}\Psi[C,t]}{\delta\sigma_{i}(\theta_{1})\,\delta\sigma_{j}(\theta_{2})}.$ (16)
Choosing a “hairpin” loop $C_{\mathrm{HP}}$ with tips at $\mathbf{r}_{1},\mathbf{r}_{2}$ and double strands between them (Fig. 1), one finds
$\Big{\langle}\omega_{i}(\mathbf{r}_{1})\,\omega_{j}(\mathbf{r}_{2})\Big{\rangle}_{v}=\left(\frac{\nu}{\mathrm{i}\,}\right)^{2}\left.\frac{\delta^{2}\Psi[C,t]}{\delta\sigma_{i}(\theta_{1})\,\delta\sigma_{j}(\theta_{2})}\right|_{C=C_{\mathrm{HP}}}.$ (17)
For the Euler ensemble this was computed exactly in Migdal (2024), yielding Mellin–Barnes integrals with ratios of zeta functions multiplied by exponentials of complex integrals of elementary functions; these are amenable to high-precision evaluation.
Figure 1: The hairpin loop $C_{\mathrm{HP}}$ with little circles at $\boldsymbol{r}_{1},\boldsymbol{r}_{2}$, standing for area derivatives, and connecting wires in between. The circulation $\Gamma_{C}[v]$ vanishes for such loop.
Thus the loop solution (Euler ensemble) gives the vorticity correlation in quadrature, with an analytically calculable spectrum of decay indices (rational and complex, the latter tied to zeros of $\zeta$). The energy spectrum, being related to vorticity correlations in Fourier space, follows similarly.
3.4 The liquid loop equation
We now derive the liquid loop equation from the Lagrangian form of the circulation dynamics. The basic object is the loop functional
$\Psi[C,t]=\Big{\langle}\exp\left(\frac{\mathrm{i}\,}{\nu}\,\Gamma_{C}[v]\right)\Big{\rangle}_{v},$ (18)
whose Lagrangian time derivative reads
$\frac{D}{Dt}\Psi[C,t]=\frac{\mathrm{i}\,}{\nu}\,\Big{\langle}\oint d\theta\ \mathbf{C}^{\prime}(\theta)\cdot\big{(}-\nu\,\nabla\times\boldsymbol{\omega}\big{)}\,\exp\left(\frac{\mathrm{i}\,}{\nu}\,\Gamma_{C}[v]\right)\Big{\rangle}_{v}.$ (19)
Using the identities of the previous subsection, we replace the gradient and vorticity by loop operators acting on the exponential:
$\nabla\;\Rightarrow\;\frac{\partial}{\partial\mathbf{C}(\theta)}=\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta-0)}-\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta+0)},$ (20)
$\hat{\omega}(\theta)\;\Rightarrow\;-\,\mathrm{i}\,\,\nu\,\frac{\delta}{\delta\boldsymbol{\sigma}(\theta)}=-\,\mathrm{i}\,\,\nu\,\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta-0)}\times\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta+0)}.$ (21)
Taking these linear operators outside the average yields a closed linear functional equation,
$\frac{D}{Dt}\Psi[C,t]=-\,\nu\,\mathcal{D}_{C}\,\Psi[C,t],$ (22)
$\mathcal{D}_{C}=\oint d\theta\ \mathbf{C}^{\prime}(\theta)\cdot\Big{(}\big{(}\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta-0)}-\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta+0)}\big{)}\times\big{(}\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta-0)}\times\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta+0)}\big{)}\Big{)}.$ (23)
3.5 From velocity dynamics to diffusion of liquid loops
Equation (22) is a diffusion equation in loop space: the advected loop executes a stochastic motion whose amplitude is linearly related to neighbouring configurations in loop space. In infinitesimal form,
$\Psi[C,t]=\big{(}I-\nu\,dt\,\mathcal{D}_{C}\big{)}\,\Psi[C,t-dt],$ (24)
and iterating produces a sum over loop histories.
As observed in earlier work, the equation becomes algebraic after a Fourier transform in loop space (“momentum loop space”). Writing
$\Psi[C,t]=\Big{\langle}\exp\left(\mathrm{i}\,\oint d\theta\ \mathbf{C}^{\prime}(\theta)\cdot\mathbf{P}(\theta)\right)\Big{\rangle}_{P},$ (25)
$\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta)}\,\Big{\langle}\exp\left(\mathrm{i}\,\oint\mathbf{C}^{\prime}\cdot\mathbf{P}\right)\Big{\rangle}_{P}=\Big{\langle}\mathrm{i}\,\,\mathbf{P}(\theta)\,\exp\left(\mathrm{i}\,\oint\mathbf{C}^{\prime}\cdot\mathbf{P}\right)\Big{\rangle}_{P},$ (26)
one finds that (22) holds provided $\mathbf{P}(\theta)$ obeys
$\partial_{t}\mathbf{P}_{\mathrm{mid}}=\nu\,\Delta\mathbf{P}\times\big{(}\Delta\mathbf{P}\times\mathbf{P}_{\mathrm{mid}}\big{)},$ (27)
$\Delta\mathbf{P}=\mathbf{P}(\theta+0)-\mathbf{P}(\theta-0),\quad\mathbf{P}_{\mathrm{mid}}=\tfrac{1}{2}\big{(}\mathbf{P}(\theta+0)+\mathbf{P}(\theta-0)\big{)}.$ (28)
This nonlinear jump condition on the unit circle resembles a Riemann–Hilbert problem (with discontinuities coupling neighbouring points rather than opposite sides of a cut). It appeared formidable in Migdal (1995) but admits an exact solution via a limiting procedure that reduces it to a Markov chain; we return to this after discussing initial data.
3.6 Initial data and the decaying regime
Formally, one could specify $\mathbf{P}(\theta,0)=\mathbf{P}_{0}(\theta)$ and reconstruct its distribution from $\Psi[C,0]$ by a functional Fourier transform. In practice, for the decaying regime of interest, the long-time statistics are attracted to a universal state independent of the initial data (Hopf’s viewpoint; see Ohkitani (2020)). The loop equation is effectively one-dimensional (the source $\mathbf{C}(\theta)$ depends on a single parameter), and its attractor can be constructed analytically.
3.7 Turbulent attractor for the liquid loop equation
The solution of (27) has power-law time dependence,
$\mathbf{P}(\theta,t)=\frac{\mathbf{F}(\theta)}{\sqrt{2\nu\,(t+t_{0})}},$ (29)
with $\mathbf{F}(\theta)$ satisfying
$\big{(}(\Delta\mathbf{F})^{2}-1\big{)}\,\mathbf{F}_{\mathrm{mid}}=(\Delta\mathbf{F}\cdot\mathbf{F}_{\mathrm{mid}})\,\Delta\mathbf{F}.$ (30)
As follows from this relation, $\mathbf{F}$ must have finite jumps at every angle; a limiting procedure yields $\mathbf{F}$ as a distribution on the circle. Requiring nontrivial vorticity $\boldsymbol{\omega}\propto\boldsymbol{P}_{\mathrm{mid}}\times\boldsymbol{\Delta}P$ excludes the trivial fixed point $\Psi[C,t]=1$ and imposes
$\Delta\mathbf{F}\cdot\mathbf{F}_{\mathrm{mid}}=0,\qquad(\Delta\mathbf{F})^{2}=1.$ (31)
Discretizing the circle at $\theta_{k}=2\pi k/N$ gives
$\mathbf{F}_{k+1}^{2}=\mathbf{F}_{k}^{2},\qquad(\mathbf{F}_{k+1}-\mathbf{F}_{k})^{2}=1.$ (32)
A particular solution is a circle in the $xy$-plane, up to a global rotation:
$\mathbf{F}_{k}=\Omega\cdot\{R\cos\alpha_{k},\ R\sin\alpha_{k},\ 0\},\quad\Omega\in SO(3),$ (33)
with
$2R^{2}\big{(}1-\cos(\alpha_{k+1}-\alpha_{k})\big{)}=1.$ (34)
Hence
$\alpha_{k+1}=\alpha_{k}+\beta\,\sigma_{k},\quad\sigma_{k}=\pm 1,\quad R=\frac{1}{2\sin(\beta/2)}.$ (35)
Periodicity $\alpha_{N}=\alpha_{0}$ yields the quantization of the angle $\beta$
$\beta=\frac{2\pi p}{q},\qquad\sum_{k=1}^{N}\sigma_{k}=r\,q,\qquad\frac{p}{q}\in\mathbb{Q},\ r\in\mathbb{Z},\ \sigma_{k}\in\mathbb{Z}_{2},$ (36)
with $N$ and $q$ even Migdal (2024). This “Euler ensemble” $\mathcal{E}$ links Euler fluid dynamics to Euler totients
$\varphi(q)=\sum_{0<p<q;\ \gcd(p,q)=1}1\qquad\text{with}\qquad\varphi(q)=q\prod_{p\mid q}\Big{(}1-\frac{1}{p}\Big{)},$ (37)
and realizes a periodic Markov chain: a random walk on the regular star polygon $\{q,p\}$. For a fixed denominator $q$, the number of admissible steps $p$ (coprime to $q$) is exactly $\varphi(q)$. Thus, when averaging over the ensemble, each family of polygons with denominator $q$ contributes with multiplicity $\varphi(q)$, and the prime decomposition of $q$ (through $\varphi$) enters the ensemble averages of the momentum–loop functional and all derived observables.
Collecting the elements, the solution of the liquid loop equation for the pure loop functional in decaying turbulence is
$\Psi[C,t]=\Big{\langle}\exp\left(\frac{\mathrm{i}\,\,\oint d\theta\ \mathbf{C}^{\prime}(\theta)\cdot\mathbf{F}(\theta)}{\sqrt{2\nu\,(t+t_{0})}}\right)\Big{\rangle}_{\mathcal{E}},$ (38a)
$\mathbf{F}(\theta)=R\,\hat{\Omega}\cdot\{\cos\alpha(\theta),\ \sin\alpha(\theta),\ 0\},\quad\hat{\Omega}\in SO(3),$ (38b)
$\alpha(\theta)=\beta\sum_{l<N\theta/(2\pi)}\sigma_{l},\qquad\sigma_{l}=\pm 1,$ (38c)
$R=\frac{1}{2\sin(\beta/2)},$ (38d)
$\beta=\frac{2\pi p}{q},\ \gcd(p,q)=1,\qquad\sum_{l=1}^{N}\sigma_{l}=r\,q,\ r\in\mathbb{Z},$ (38e)
$N\bmod 2=0.$ (38f)
This is the discretized loop solution: the loop is approximated by an $N$-gon with sides $\Delta\mathbf{C}_{k}=\mathbf{C}_{k}-\mathbf{C}_{k-1}=O(1/N)$. The local limit $N\to\infty$ is taken together with $\nu\to 0$ at fixed $\tilde{\nu}=\nu N^{2}$.
3.8 Back to the static loop equation
We now map the Euler-ensemble solution of the liquid loop equation back to the static loop equation by prescribing a loop motion in momentum space that reproduces the missing $\mathbf{v}\times\boldsymbol{\omega}\,$ term. Consider
$\Psi[C,t]=\Big{\langle}\exp\left(\mathrm{i}\,\oint d\theta\ \mathbf{C}^{\prime}(\theta)\cdot\mathbf{P}(\theta)\right)\Big{\rangle}_{\mathcal{E}},$ (39)
$\partial_{t}\mathbf{C}(\theta)=-\,\mathbf{v}_{P}(\theta),$ (40)
$\frac{D}{Dt}\Psi[C,t]=\partial_{t}\Psi[C,t]-\Big{\langle}\mathrm{i}\,\oint d\theta\ \partial_{\theta}\mathbf{v}_{P}(\theta)\cdot\mathbf{P}(\theta)\,\exp\left(\mathrm{i}\,\oint\mathbf{C}^{\prime}\cdot\mathbf{P}\right)\Big{\rangle}_{\mathcal{E}}.$ (41)
Integrating by parts and using $\mathrm{i}\,\,\partial_{\theta}\mathbf{P}\Rightarrow-\,\delta/\delta\mathbf{C}$ on the exponential yields the desired static advection operator. With the dictionary (20),
$\nabla\Rightarrow\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta-0)}-\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta+0)}\Rightarrow-\,\mathrm{i}\,\,\Delta\mathbf{P},$ (42)
$\boldsymbol{\omega}\Rightarrow\mathrm{i}\,\nu\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta-0)}\times\frac{\delta}{\delta\mathbf{C}^{\prime}(\theta+0)}\Rightarrow\mathrm{i}\,\nu\mathbf{P}_{\mathrm{mid}}\times\Delta\mathbf{P},$ (43)
so that
$\mathbf{v}_{P}=\nu\frac{\Delta\mathbf{P}\times(\mathbf{P}_{\mathrm{mid}}\times\Delta\mathbf{P})}{(\Delta\mathbf{P})^{2}}\quad\stackrel{{\scriptstyle\mathcal{E}}}{{\Longrightarrow}}\quad\mathbf{v}_{P}=\nu\mathbf{P}_{\mathrm{mid}}.$ (44)
The extra term in momentum space is then
$\mathrm{i}\,\oint d\theta\ \partial_{\theta}\mathbf{v}_{P}\cdot\mathbf{P}=-\,\mathrm{i}\,\oint d\theta\ \mathbf{v}_{P}\cdot\partial_{\theta}\mathbf{P}=\mathrm{i}\,\nu\oint d\theta\ \mathbf{P}_{\mathrm{mid}}\cdot\partial_{\theta}\mathbf{P},$ (45)
which vanishes: the smooth part $\mathbf{P}_{\mathrm{mid}}$ of $\mathbf{P}$ gives a total derivative over the closed loop, and the jump contributions cancel by the Euler-ensemble constraints. Equivalently, solving for the moving momentum loop
$\mathbf{C}(\theta,t)=\mathbf{C}_{0}(\theta)+\big{(}\sqrt{2\nu(t+t_{0})}-\sqrt{2\nu t_{0}}\big{)}\,\mathbf{F}_{\mathrm{mid}}(\theta)$ (46)
one finds that the time-dependent term cancels as in the liquid equation. Thus the Euler ensemble solves both the liquid and static loop equations.
3.9 Euler-ensemble predictions and validation
The rigorous analysis of the diffusion part and the discrete-to-continuum limit is in preparation Brue & DeLellis (2025). A physics-level proof using a discrete Euler ensemble and verifying the Navier-Stokes loop equation up to $O(1/N)$ corrections is given in Migdal (2025a), and now we extended that analysis with the equivalence of the static and liquid formulations. Using (17), the Euler-ensemble predictions were compared in Migdal (2024, 2025a) with both laboratory and DNS data. For example, turbulent kinetic energy decays as $t^{-5/4}$ in the Euler ensemble, while the classic grid-turbulence measurements Comte-Bellot & Corrsin (1966) found $t^{-1.25\pm 0.01}$; the Kolmogorov–Saffman model would give $t^{-1.20}$, outside the experimental error bar. Likewise, the effective index (log–log derivative) of the second-order velocity increment as a function of $\log|\mathbf{r}|/\sqrt{t}$ follows the predicted nonlinear curve and differs markedly from the constant K41 value $2/3$ and from constant-index multifractal variants.
In what follows we take the velocity solution as given and apply its statistics (the Euler ensemble) to the passive-scalar field.
4 Passive scalar dynamics in loop space
In the conventional approach Falkovich et al. (2001), the advection–diffusion equation is coupled to an unknown velocity field; progress typically comes from Gaussian surrogates or ad hoc closures of the moment hierarchy. In the loop formulation, the problem reduces to a closed linear equation for the loop functional, which we solve exactly—no surrogates, no closures.
4.1 Circulation and temperature joint distribution
Consider the passive scalar (temperature) evaluated at a loop point $\boldsymbol{r}_{1}=\boldsymbol{C}(\theta_{1})$. The advection–diffusion equation can be written in terms of loop operators as
$\partial_{t}T(\boldsymbol{C}(\theta_{1}),t)=\left(-\boldsymbol{v}(\boldsymbol{C}(\theta_{1}))\cdot\partial_{\boldsymbol{C}(\theta_{1})}+D\left(\partial_{\boldsymbol{C}(\theta_{1})}\right)^{2}\right)T(\boldsymbol{C}(\theta_{1}),t)$ (47)
Note that the gradient $\partial_{\boldsymbol{C}(\theta_{1})}$ represents the singular limit in which only a single point on the loop is shifted; we return to this point in the next subsection.
Because the scalar does not feed back on the flow, the Euler ensemble (as established in Migdal (2023, 2024) and reviewed above) continues to describe the velocity statistics. Diffusion affects only the scalar statistics, so we must solve the advection–diffusion problem in loop space given by (47).
To this end, consider the mixed average
$T_{C}(\theta_{1},t)=\left\langle T\left(\boldsymbol{C}(\theta_{1}),t\right)\exp\left(\frac{\mathrm{i}\,\Gamma_{C}[v]}{\nu}\right)\right\rangle_{v}$ (48)
which correlates the temperature at $\boldsymbol{C}(\theta_{1})$ with the circulation around $C$. Differentiating in time produces contributions both from (47) and from the Navier-Stokes dynamics:
$\partial_{t}T_{C}(\theta_{1},t)=\left\langle\left(-\boldsymbol{v}(\boldsymbol{C}(\theta_{1}))\cdot\partial_{\boldsymbol{C}(\theta_{1})}+D\partial_{\boldsymbol{C}(\theta_{1})}^{2}\right)T\left(\boldsymbol{C}(\theta_{1}),t\right)\exp\left(\frac{\mathrm{i}\,\Gamma_{C}[v]}{\nu}\right)\right\rangle_{v}$
$+\frac{\mathrm{i}\,}{\nu}\left\langle T\left(\boldsymbol{C}(\theta_{1}),t\right)\oint d\theta\boldsymbol{C}^{\prime}(\theta)\left(-\nu\boldsymbol{\nabla}\times\boldsymbol{\omega}-\boldsymbol{\omega}\times\boldsymbol{v}\right)(\boldsymbol{C}(\theta))\exp\left(\frac{\mathrm{i}\,\Gamma_{C}[v]}{\nu}\right)\right\rangle_{v}$ (49)
4.2 The loop equation for the passive scalar
We now rewrite (4.1) as a loop equation for $T_{C}$:
$\partial_{t}T_{C}(\theta_{1},t)=\left\langle\left(-\hat{v}(\theta_{1})\cdot\partial_{\boldsymbol{C}(\theta_{1})}+D\partial_{\boldsymbol{C}(\theta_{1})}^{2}\right)T\left(\boldsymbol{C}(\theta_{1}),t\right)\exp\left(\frac{\mathrm{i}\,\Gamma_{C}[v]}{\nu}\right)\right\rangle_{v}$
$+\frac{\mathrm{i}\,}{\nu}\left\langle T\left(\boldsymbol{C}(\theta_{1}),t\right)\oint d\theta\boldsymbol{C}^{\prime}(\theta)\cdot\left(-\nu\partial_{\boldsymbol{C}(\theta)}\times\hat{\omega}(\theta)+\hat{\omega}(\theta)\times\hat{v}(\theta)\right)\exp\left(\frac{\mathrm{i}\,\Gamma_{C}[v]}{\nu}\right)\right\rangle_{v};$ (50)
$\hat{\omega}(\theta)=-\mathrm{i}\,\nu\frac{\delta}{\delta\boldsymbol{\sigma}(\theta)};$ (51)
$\hat{v}(\theta)=\frac{-1}{\partial_{\boldsymbol{C}(\theta)}^{2}}\partial_{\boldsymbol{C}(\theta)}\times\hat{\omega}(\theta);$ (52)
Taking the loop operators outside the average $\left\langle\cdot\right\rangle_{v}$ would, in principle, generate cross terms when $\partial_{\boldsymbol{C}(\theta_{1})}$ acts on $\exp\left(\frac{\mathrm{i}\,\Gamma_{C}[v]}{\nu}\right)$ and when $\frac{\delta}{\delta\boldsymbol{\sigma}(\theta)}$ acts on $T(\boldsymbol{C}(\theta_{1}),t)$. All such cross terms vanish by the identities (12) and (8). Hence we obtain the closed loop equation
$\partial_{t}T_{C}(\theta_{1},t)=\left(-\hat{v}(\theta_{1})\cdot\partial_{\boldsymbol{C}(\theta_{1})}+D\partial_{\boldsymbol{C}(\theta_{1})}^{2}\right)T_{C}(\theta_{1},t)$
$+\frac{\mathrm{i}\,}{\nu}\oint d\theta\boldsymbol{C}^{\prime}(\theta)\cdot\left(-\nu\partial_{\boldsymbol{C}(\theta)}\times\hat{\omega}(\theta)-\hat{\omega}(\theta)\times\hat{v}(\theta)\right)T_{C}(\theta_{1},t);$ (53)
4.3 Solution of the passive scalar loop equation
The Euler-ensemble average for the passive scalar has the form
$T_{C}(\theta_{1},t)=\left\langle\tau(\boldsymbol{C}(\theta_{1}),t)\exp\left(\frac{\mathrm{i}\,\oint d\theta\boldsymbol{C}^{\prime}(\theta)\cdot\boldsymbol{F}(\theta)}{\sqrt{2\nu(t+t_{0})}}\right)\right\rangle_{\mathcal{E}}$ (54)
where $\tau$ is determined by
$\frac{\partial\tau(\boldsymbol{C}(\theta_{1}),t)}{\partial t}\exp\left(\frac{\mathrm{i}\,\oint d\theta\boldsymbol{C}^{\prime}(\theta)\cdot\boldsymbol{F}(\theta)}{\sqrt{2\nu(t+t_{0})}}\right)=\hat{H}(\theta_{1})\tau(\boldsymbol{C}(\theta_{1}),t)\exp\left(\frac{\mathrm{i}\,\oint d\theta\boldsymbol{C}^{\prime}(\theta)\cdot\boldsymbol{F}(\theta)}{\sqrt{2\nu(t+t_{0})}}\right);$ (55)
$\hat{H}(\theta)=-\tilde{v}(\theta,t)\cdot\partial_{\boldsymbol{C}(\theta)}+D\partial_{\boldsymbol{C}(\theta)}^{2}$ (56)
$\tilde{v}(\theta_{1},t)=\frac{1}{\Delta\boldsymbol{F}^{2}(\theta_{1})}\Delta\boldsymbol{F}(\theta_{1})\times(\boldsymbol{F}(\theta_{1})\times\Delta\boldsymbol{F}(\theta_{1}))\frac{\nu}{\sqrt{2\nu(t+t_{0})}};$ (57)
and the remaining terms in (4.2) cancel by the loop equation for $\Psi$,
$\partial_{t}\Psi(C,t)=\frac{\mathrm{i}\,}{\nu}\oint d\theta\boldsymbol{C}^{\prime}(\theta)\cdot\hat{L}(\theta)\Psi(C,t);$ (58)
$\hat{L}(\theta)=-\nu\partial_{\boldsymbol{C}(\theta)}\times\hat{\omega}(\theta)-\hat{\omega}(\theta)\times\hat{v}(\theta)$ (59)
Using (11) and the cancellations noted above, we obtain a linear evolution for $\tau$,
$\frac{\partial\tau(\boldsymbol{C}(\theta_{1}),t)}{\partial t}=\hat{Q}(\theta_{1},t)\tau(\boldsymbol{C}(\theta_{1}),t);$ (60)
$\hat{Q}(\theta,t)=\left(\frac{-2\mathrm{i}\,D\Delta\boldsymbol{F}(\theta)}{\sqrt{2\nu(t+t_{0})}}-\tilde{v}(\theta,t)\right)\cdot\partial_{\boldsymbol{C}(\theta)}+D\partial_{\boldsymbol{C}(\theta)}^{2}-D\frac{(\Delta\boldsymbol{F}(\theta))^{2}}{2\nu(t+t_{0})};$ (61)
$\Delta\boldsymbol{F}(\theta_{k})=\boldsymbol{F}(\theta_{k})-\boldsymbol{F}(\theta_{k-1});$ (62)
With $|\Delta\boldsymbol{F}(\theta_{k})|^{2}=1$ one finds
$\tilde{v}(\theta_{k},t)=\frac{\nu(\boldsymbol{F}(\theta_{k})+\boldsymbol{F}(\theta_{k-1}))}{2\sqrt{2\nu(t+t_{0})}};$ (63)
and Fourier representation gives
$\tau(\boldsymbol{r},t)=\int\frac{d^{3}q}{(2\pi)^{3}}\exp\left(\mathrm{i}\,\boldsymbol{q}\cdot\boldsymbol{r}\right)\eta(\boldsymbol{q},t);$ (64)
$\eta(\boldsymbol{q},t)=W(\boldsymbol{q})\left(\frac{t+t_{0}}{t_{0}}\right)^{-\frac{1}{2\mathit{Sc}}}\exp\left(-\frac{\boldsymbol{q}^{2}\nu t}{\mathit{Sc}}-\mathrm{i}\,\left(\sqrt{2\nu(t+t_{0})}-\sqrt{2\nu t_{0}}\right)\boldsymbol{q}\cdot\boldsymbol{G}_{k}\right);$ (65)
$\boldsymbol{G}_{k}=\frac{\boldsymbol{F}(\theta_{k})+\boldsymbol{F}(\theta_{k-1})}{2}+2\mathrm{i}\,\frac{\boldsymbol{F}(\theta_{k})-\boldsymbol{F}(\theta_{k-1})}{\mathit{Sc}}$ (66)
Here $\mathit{Sc}=\nu/D$ is the Schmidt number, the sole dimensionless parameter for the scalar. We also relabel the picked angle as $\theta_{k}$ to emphasize that $\theta_{k}=2\pi k/N$ tends to a continuum as $N\to\infty$.
The weight $W(\boldsymbol{q})$ reflects the initial scalar distribution:
$W(\boldsymbol{q})=\eta(\boldsymbol{q},0)=\int d^{3}r\exp\left(-\mathrm{i}\,\boldsymbol{q}\cdot\boldsymbol{r}\right)T(\boldsymbol{r},0)$ (67)
The Fourier quantity $\eta(\boldsymbol{q},t)$ is complex because the temperature is averaged with the phase factor $\exp\left(\mathrm{i}\,\Gamma_{C}[v]/\nu\right)$.
The mean temperature is, of course, real and positive. To extract it, contract the loop to a point: set $\boldsymbol{r}=\boldsymbol{C}(\theta_{1})$ and $\boldsymbol{C}^{\prime}(\theta)=0$. Then $\Gamma_{C}[v]=0$, and the original average (48) reduces to the scalar mean at a point. On the right of (54), setting $\boldsymbol{C}^{\prime}=0$ eliminates the momentum-loop factor. Thus
$\left\langle T(\boldsymbol{r},t)\right\rangle_{v}=\left\langle\tau(\boldsymbol{C}(\theta_{1}),t)\right\rangle_{\mathcal{E}}=\left(\frac{t+t_{0}}{t_{0}}\right)^{-\frac{1}{2\mathit{Sc}}}$
$\int\frac{d^{3}q}{(2\pi)^{3}}W(\boldsymbol{q})\exp\left(\mathrm{i}\,\boldsymbol{q}\cdot\boldsymbol{r}\right)\left\langle\exp\left(-\frac{\boldsymbol{q}^{2}\nu t}{\mathit{Sc}}-\mathrm{i}\,\left(\sqrt{2\nu(t+t_{0})}-\sqrt{2\nu t_{0}}\right)\boldsymbol{q}\cdot\boldsymbol{G}_{k}\right)\right\rangle_{\mathcal{E}};$ (68)
4.4 The turbulent limit. Harmony behind chaos
As shown in Migdal (2023, 2024), the turbulent limit combines infinitesimal viscosity with large loop amplitude:
$\nu\to\frac{\tilde{\nu}}{N^{2}}=O(1/N^{2});$ (69)
$\rho=R/N=O(1);$ (70)
$\boldsymbol{F}(\theta)\to N\rho\left\{\cos\alpha,\sin\alpha,0\right\};$ (71)
$|\Delta\boldsymbol{F}(\theta)|=1;$ (72)
In this limit the vector $\sqrt{\nu}\boldsymbol{G}_{k}$ tends to a finite real vector,
[…]
and the statistics of $\rho=\big{(}2N\sin(\pi p/q)\big{)}^{-1}$ for coprime $p,q$ with $q\sim N\to\infty$ are (as in Migdal (2024)) a discontinuous, piecewise power-like distribution:
$f(\rho)=\left(1-\frac{\pi^{2}}{21600\zeta(5)}\right)\delta(\rho)+\frac{2\pi^{3}}{3}\rho^{4}\Phi\left(\left\lfloor\frac{1}{2\pi\rho}\right\rfloor\right);$ (74)
with $\Phi$ the totient summatory function:
$\Phi(q)=\sum_{n=1}^{q}\varphi(n)$ (75)
For a point-source initial condition,
$T(\boldsymbol{r},0)=T_{0}\delta(\boldsymbol{r})$ (76)
we find from (64)
$\left\langle T(\boldsymbol{r},t)\right\rangle_{v}=T_{0}(1+t/t_{0})^{-\frac{1}{2\mathit{Sc}}}G(r,t);$ (77)
$G(r,t)=\left\langle\delta\left(\boldsymbol{r}-\rho\left(\sqrt{2\tilde{\nu}(t+t_{0})}-\sqrt{2\tilde{\nu}t_{0}}\right)\hat{\Omega}\cdot\left\{\cos\alpha_{k},\sin\alpha_{k},0\right\}\right)\right\rangle_{\rho,\hat{\Omega}}$ (78)
The one-dimensional delta in (74) produces a three-dimensional delta in $G$; the remaining term yields a finite structure. Averaging over $\hat{\Omega}\in\mathbb{O}(3)$ gives
$G(r,t)=\left(1-\frac{\pi^{2}}{21600\zeta(5)}\right)\delta(\boldsymbol{r})+\frac{\pi^{2}}{6r^{2}}\int d\rho\rho^{4}\delta\left(r-2\pi\rho L(t)\right)\Phi\left(\left\lfloor\frac{1}{2\pi\rho}\right\rfloor\right);$ (79)
$L(t)=\frac{\sqrt{2\tilde{\nu}(t+t_{0})}-\sqrt{2\tilde{\nu}t_{0}}}{2\pi};$ (80)
and after integrating over $\rho$,
$\left\langle T(\boldsymbol{r},t)\right\rangle_{v}=(1+t/t_{0})^{-\frac{1}{2\mathit{Sc}}}\left[\left(1-\frac{\pi^{2}}{21600\zeta(5)}\right)\delta(\boldsymbol{r})+\frac{r^{2}\Phi\left(\left\lfloor\frac{L(t)}{r}\right\rfloor\right)}{768\pi^{5}L(t)^{5}}\right];$ (81)
This is a decaying prefactor multiplying a rescaled delta function plus a second, decaying contribution with a discontinuous dependence on the scaling variable $r/L(t)$ (Fig. 2). The support lies at $r\leq L(t)$, so the shells contract towards the origin as $t\downarrow 0$.
Thus a point source produces infinitely many concentric spherical shells, $\frac{L(t)}{n}<r<\frac{L(t)}{n-1}$, accumulating at the origin. The profile is parabolic within each shell, with jumps $\propto\varphi(n)/n^{2}$ at successive boundaries. Since these jumps decrease with $n$, the temperature approaches a finite limit at the origin:
$\lim_{n\to\infty}\frac{\Phi\left(n\right)}{n^{2}}=\frac{3}{\pi^{2}}$ (82)
The discontinuities reflect the absence of a continuum limit for Euler’s totient; e.g.,
$\varphi(7919)=7818;$ (83)
$\varphi(7920)=1920;$ (84)
although statistically, for large $n$,
$\varphi(n)\sim\textit{ const }{}\frac{n}{\log\log n}\quad\text{for almost all }n$ (85)
Observing such a shell structure directly in DNS would require very large grids; in physical flows it may be smoothed by weak dissipation or forcing. The stability of the shell pattern is an open problem that merits further study.
Figure 2: Log–log plot of the universal function $\xi^{2}\Phi\left(\left\lfloor\frac{1}{\xi}\right\rfloor\right)$ where $\xi=\frac{2\pi r}{\left(\sqrt{2\tilde{\nu}(t+t_{0})}-\sqrt{2\tilde{\nu}t_{0}}\right)}$.
For spectral properties, take $W(\boldsymbol{q}^{\prime})=\delta(\boldsymbol{q}-\boldsymbol{q}^{\prime})$ (a single Fourier mode):
$\left\langle\tilde{T}(\boldsymbol{q},t)\right\rangle_{v}=\left(1+\frac{t}{t_{0}}\right)^{-\frac{1}{2\mathit{Sc}}}U\left(|\boldsymbol{q}|\left(\sqrt{2\tilde{\nu}(t+t_{0})}-\sqrt{2\tilde{\nu}t_{0}}\right)\right);$ (86)
$U(\kappa)=\left\langle\exp\left(\mathrm{i}\,\rho\boldsymbol{\kappa}\cdot\hat{\Omega}\cdot\left\{\cos\alpha_{k},\sin\alpha_{k},0\right\}\right)\right\rangle_{\rho,\hat{\Omega}}=\left\langle\frac{\sin(\rho\kappa)}{\rho\kappa}\right\rangle_{\rho}$ (87)
Using (74) and the Taylor expansion of $\sin(\rho\kappa)/(\rho\kappa)$ yields an absolutely convergent series
$U(\kappa)=1+\sum_{l=1}^{\infty}\frac{(l+1)(2\pi)^{-2l-3}\left(-\kappa^{2}\right)^{l}\zeta(2(l+2))}{3(2l+5)\zeta(2l+5)\Gamma(2l+3)}$ (88)
We have evaluated this series numerically in Mathematica ®  with high precision: $U(\kappa)$ starts at $1$, decreases slightly, and then oscillates about $0.99986$ (Fig. 3).
Figure 3: The plot of universal function $U(\kappa)$ in (88) describing the amplitude of passive scalar plane wave with wavevector $\boldsymbol{q}$ at the time $t$ as a function of scaling variable $\kappa=|\boldsymbol{q}|\left(\sqrt{2\tilde{\nu}(t+t_{0})}-\sqrt{2\tilde{\nu}t_{0}}\right)$
Why discrete shells survive at fixed Sc
We now explain, in standard scaling terms, why discrete shells appear and persist at fixed $\mathit{Sc}$. In the decaying, high–Re limit we take $\nu\!\to\!0$ and $D\!\to\!0$ at fixed $\mathrm{Sc}=\nu/D$. Kelvin cancellation leaves a closed loop equation whose nonlinear balance sustains finite vorticity and a finite advective speed (of order unity), whereas the diffusive speed scales as $v_{D}\sim\sqrt{D/t}\ll 1$. Thus advection dominates mixing, with diffusion acting as a secondary smoothener.
In this regime the scalar is transported ballistically to radii $r=\rho\,L(t)$, with $L(t)\sim\sqrt{t}$ and a dimensionless transport ratio $\rho$ that does not vanish as $\nu\!\to\!0$. The admissible values of $\rho$ form a discrete set fixed by a closure/periodicity constraint; for each denominator $q$ the number of admissible numerators (coprime to $q$) is $\varphi(q)$ (Euler’s totient). Because the advection–diffusion equation is linear in $T$, the ensemble superposes these discrete transports, producing concentric shells at $r=L(t)/n$ with parabolic profiles inside each shell and jumps proportional to $\varphi(n)/n^{2}$ across shells. Any finite $D$ rounds the discontinuities but leaves the discrete radii and the jump law intact at leading order. In Fourier variables this structure appears as a universal function of $|\mathbf{q}|\,\big{(}\sqrt{t+t_{0}}-\sqrt{t_{0}}\big{)}$ with small, parameter–free oscillations.
A constructive realization of the admissible transport ratios is given in the momentum–loop formulation (see (37)), but it is not needed for the scaling argument above.
5 Discussion. Myths and reality of decaying turbulence
Figure 4: The turbulent kinetic energy decaying with time in the $4K$ grid simulation with initial Gaussian velocity field distributed according to the K41 spectrum $k^{-5/3}$. After a few hundred time steps, the statistical equilibrium was reached, and energy started decaying in agreement with our $t^{-5/4}$ law. There are other parameters of that stage of decay, with distribution matching our theory, such as the length scale $L(t)\sim\sqrt{t}$ and effective index of the second moment of velocity field which is a nontrivial function of scaling variable $\log(t/\sqrt{t})$ (see Fig. 5).
Figure 5: Verification of the Euler-ensemble theory in DNS. Upper left: parabolic relation between time and the effective length $L(t)$ for the $4K$ grid. Upper right: leading law $E\propto L(t)^{-5/2}$ (blue) and the full theoretical curve (green) vs. DNS (red) for $4K$. Lower right: same comparison for a smaller $1K$ grid. Lower left: effective index (DNS with error bars, red), theory (green), and the K41 value (blue dotted).
The loop equations as a route to decaying turbulence represent a departure from customary practice, and they naturally meet resistance from the community. The objection is often framed as follows: rather than resolving the myriad vortical structures and their interactions in physical space, the loop approach reduces the problem to the statistics of a stochastic process on regular star polygons (Fig. 6). Can such a one–dimensional description be truly equivalent to three–dimensional vortex dynamics?
A related concern is that any putative solution must address the familiar themes of the classical picture: energy cascade, vortical hierarchies, multifractal scaling, and intermittency. Where do these ideas sit in the loop description?
Figure 6: Regular star polygons, describing the Euler ensemble. There are also random Ising variables $\sigma=\pm 1$ sitting at vertices and telling the direction of the random walk from this vertex to the next or the previous one.
Our position is simple. Some widely used assumptions and scaling templates do not apply to the decaying turbulence regime; others are modified or generalized in the loop picture. Below we explain how the main questions are answered within the theory.
Do we see any star polygons?
The polygons are not meant to be seen in a flow visualization; they are a convenient basis in loop space. Their role is analogous to unobservable but predictive degrees of freedom in other areas of physics (such as quark confinement): they lead to quantitative, testable predictions, such as the decay of turbulent kinetic energy as $t^{-\nicefrac{{5}}{{4}}}$.
Where did all this complexity go?
The key is duality: the same physics can admit complementary mathematical descriptions. In one representation the dynamics look intricate (waves and eddies); in another they are encoded by a simpler stochastic object. Just as strong coupling in one theory can map to weak coupling in a dual theory, the turbulent limit of the velocity dynamics maps to a WKB–like limit of the Euler ensemble, which is exactly solvable at infinite Reynolds number. In this sense, the “complexity” is not lost; it is recast. Here it appears in the arithmetic structure (prime factors via Euler totients and zeta functions) rather than in a cascade hierarchy.
The myth of Kolmogorov scaling
With respect to the decaying regime, the classical K41 template is not supported by data at the level often assumed. Kolmogorov’s original constant–dissipation ansatz was a model assumption, and even he (with Obukhov) introduced fluctuations (log–normal) soon after. Modern experiments and DNS Yakhot & Zakharov (1993); Yakhot & Sreenivasan (2004); Sreenivasan & Yakhot (2021) show significant, systematic deviations from K41 (and from its log–normal variant). In decaying turbulence, experiments Panickacheril John et al. (2022) deviate even more strongly, despite attempts to fit or trim the data; for example, the energy spectrum departs from $k^{-\nicefrac{{5}}{{3}}}$ over many decades, and recent measurements Küchler et al. (2023) report clear departures of the effective index of $\left\langle\Delta v^{2}\right\rangle(r)$ from the K41 value $2/3$.
Multifractals and the myth of conformal symmetry
Motivated by analogies with critical phenomena, multifractal models posit power laws with anomalous exponents in an inertial range and can organize deviations from K41 in forced turbulence Parisi & Frisch (1985); Yakhot & Sreenivasan (2004); Sreenivasan & Yakhot (2021). In the incompressible turbulence, however, there is no theoretical basis for conformal symmetry: the conservation constraints of a CFT would assign equal dimensions to $\boldsymbol{v},\boldsymbol{\omega}$ which is incompatible with $\boldsymbol{\omega}=\nabla\times\mathbf{v}$. More importantly, the decaying spectra are not straight lines on log–log plots; they bend in a universal way (cf. Fig. 2 (top) of Küchler et al. (2023)), indicating that single power laws—whether K41 or multifractal—are insufficient. The loop theory predicts nonlinear, universal functions of $\log k\sqrt{t}$ across several decades, consistent with the observed curvature.
Our generalization of multifractal laws
While not conformally invariant, the loop theory retains a key “spectral” feature: the Mellin transform of the vorticity correlation is meromorphic in the Mellin parameter, yielding an infinite discrete set of complex scaling indices,
indexes of velocity correlation
$\begin{array}[]{c|c}\text{Index}&\text{Condition}\\ \hline\cr-1&\\ 0&\\ \hline\cr 2n&n\in\mathbb{Z},n\geq 1\\ 5/2&\\ 11/2&\\ \frac{15+4n}{2}&n\in\mathbb{Z},n\geq 0\\ 7\pm\imath t_{n}&n\in\mathbb{Z}\end{array}$ (97)
where $\frac{1}{2}\pm\imath t_{n}$ are zeros of $\zeta$. This extends the multifractal idea (many exponents) to a fully discrete arithmetic spectrum. Recent DNS Sreenivasan (2023) report slopes consistent with these predictions, unlike the classical scaling templates of Kolmogorov or Saffman.
The $4/5$ law and the cascade narrative
The $4/5$ law is a celebrated and exact statement about third–order moments $\left\langle vvv\right\rangle$. In the present context it does not constrain the rotational part $\left\langle vv\omega\right\rangle$ and it does not enforce a constant spectral flux. Indeed, the linear-in-$\mathbf{r}$ term transforms to gradients of delta functions in $\mathbf{k}$–space rather than to a scale–independent flux across wavenumbers. Within the loop theory, the energy spectrum decays fast enough (in the turbulent limit) that the Kolmogorov flux vanishes, consistent with classical bounds Sulem & Frisch (1975). The dissipative anomaly reconciles finite dissipation with this vanishing flux: our exact expressions for dissipation and spectra Migdal (2024) are compatible with the anomaly but not with a cascade picture in decaying turbulence.
The myth of Heisenberg scaling
Heisenberg’s $k^{-7}$ proposal for a dissipative subrange, taken up by Chandrasekhar, was a model of its time Heisenberg (1948); Chandrasekhar (1949). As Sreenivasan recounts Sreenivasan (2019a), von Neumann already noted in 1949 the lack of experimental support. In our Mellin spectrum there is no pole between 13/2-13/2 and $8\pm\mathrm{i}t_{n}$, and the effective index approaches 7/2-7/2 rather than 7-7. In other words, the nontrivial dynamics persist across what used to be separated as “inertial” and “dissipative” ranges. The resulting universal decaying spectrum spans several decades, and the effective index of $\left\langle\Delta v^{2}\right\rangle(r)$ varies smoothly from $2$ down to 0, with no plateau at $2/3$. These predictions match DNS and experiments across the entire turbulent range without adjustable dimensionless parameters.
Forced versus spontaneous stochasticity
Random forcing is a modeling device to maintain stationarity. Decaying turbulence exhibits spontaneous stochasticity: over time, the measure spreads from near–deterministic initial data to a broad distribution over solutions. Our loop equation does not attempt to model the bifurcation pathway; it characterizes the attractor reached in the turbulent limit. On that attractor, the physics is qualitatively different from the cascade narrative: the dynamics are organized by a loop–space diffusion and an arithmetic spectrum.
Is forced turbulence universal?
With forcing kept finite, universality can depend on how the forcing is implemented. In particular, the large-wavelength forcing of the potential velocity component with long–range correlations $\left\langle vv\right\rangle(r)\sim r^{\alpha},\alpha\approx 0.67$ can seed non-universal large–scale structure, so that the flow reflects the driving more than the fluid dynamics. By contrast, the loop equation isolates the vorticity sector, whose correlations decay with distance and are less sensitive to boundaries and driving. This separation helps identify a universal decaying–turbulence attractor.
Grid turbulence and DNS
A classic setting is grid turbulence (Fig. 7).
Figure 7: Typical setup of the grid turbulence experiment. The flow enters from the left, passing through the oscillating grid, creating vorticity. These vortexes interact, exchange energy and eventually reach some statistical equilibrium, with energy constantly dissipated in the bulk, while being resupplied from the flow through the grid. In the still Galilean frame, this is a steady process, with energy coming from the left boundary and dissipated in the bulk. The local kinetic energy decays with the distance $z$ from the grid as $z^{-1.25}$, but stays constant with time. In the frame moving with the mean velocity of the flow, this is decaying turbulence, with energy decaying with time as $(vt)^{-1.25}$.
The landmark experiments of Comte-Bellot & Corrsin (1966) measured a decay exponent $1.25\pm 0.01$, close to $5/4$ and inconsistent with the Kolmogorov–Saffman value $6/5$. More recently, DNS on $4\mathrm{K}$ grids with periodic boundaries and K41-type initial spectra (no forcing) were performed by Sreenivasan and collaborators. The simulations relaxed to a decaying state in which the energy decays as $t^{-5/4}$, the integral scale obeys $L(t)\sim t^{1/2}$, and the effective index of the second moment follows the predicted nonlinear curve (Fig. 4, 5).
Summary. In the decaying turbulence, single–power scaling laws and cascade arguments do not describe the observed behavior. The loop equation replaces these templates with a microscopic, parameter–free description: universal nonlinear spectral shapes, an arithmetic spectrum of indices, and a diffusion in loop space that organizes the statistics. The evidence from classic experiments and modern DNS supports this picture across the turbulent range.
6 Conclusions
We have presented an exact analytic solution for the one–point scalar distribution from a point source in decaying turbulence, in the joint limit $\nu\to 0$, $D\to 0$ at fixed Schmidt number $\mathit{Sc}=\nu/D$. The background velocity statistics are those of the Euler ensemble. In this regime the scalar is transported ballistically by the advective field while molecular smoothing is subleading; the mean field forms sharply defined, compactly supported spherical shells whose outer radius grows as $L(t)\sim\sqrt{t}$.
Main implications.
  • Closed scalar loop equation and exact solution. The loop formulation yields a closed linear equation for the scalar loop functional and its exact solution at fixed $\mathit{Sc}$. In real space, a point source produces a hierarchy of concentric shells at discrete radii $r=L(t)/n$, each with a piecewise–parabolic profile; the jump at the $n$th shell is proportional to $\varphi(n)/n^{2}$ (totient multiplicities). These structures are the geometric skeleton of scalar transport in the decaying limit; any finite $D$ rounds the discontinuities but leaves the discrete radii at leading order.
  • Liquid vs static loop equivalence; dominance of advection. The Euler ensemble solves the liquid (Lagrangian) loop equation; on this solution the advection term in the static (Eulerian) loop equation vanishes, so both equations hold. A scaling argument at fixed $\mathit{Sc}$ explains why the advective phase stays $O(1)$ while the diffusive factor is small, which is why the shells persist in the limit.
  • Fourier signature. For a monochromatic initial condition the solution reduces to a universal, parameter–free response $U(\kappa)$ with an absolutely convergent series in $\kappa=|\mathbf{q}|\,(\sqrt{t+t_{0}}-\sqrt{t_{0}})$. The overall amplitude decays as $(1+t/t_{0})^{-1/(2\mathit{Sc})}$; $U(\kappa)$ exhibits small, predictable oscillations. This is a clean DNS observable that does not require resolving shells in real space.
  • Consistency with decaying–velocity statistics. The scalar results rest on the Euler ensemble, whose predictions for velocity decay and scales agree with classic experiments and modern DNS: turbulent kinetic energy $E(t)\sim t^{-5/4}$, $L(t)\sim t^{1/2}$, and a smoothly varying effective index for $\left\langle\Delta v^{2}\right\rangle(r)$ across the turbulent range (no plateau at $2/3$). No adjustable dimensionless parameters are used.
  • Verification strategy. Direct observation of shells may be challenging in finite–resolution flows, but the Fourier–space decay law and $U(\kappa)$ provide a robust test. In addition, the predicted shell radii $r=L(t)/n$ and the jump law $\propto\varphi(n)/n^{2}$ offer real–space diagnostics if resolution permits.
  • Scope and regimes. The theory applies to unforced, decaying turbulence in the whole space, in the high–Reynolds, vanishing–dissipation limit at fixed $\mathit{Sc}$. Weak forcing or finite diffusivity smooth the discontinuities; the discrete radii survive in the preasymptotic regime. Uniqueness and stability of the attractor are interesting mathematical questions left open.
  • Early–time behaviour in DNS. For $t\ll t_{0}$, the factor $(1+t/t_{0})^{-1/(2\mathit{Sc})}$ is close to $\exp\left(-\frac{t}{2t_{0}\mathit{Sc}}\right)$, so long runs are beneficial for separating the algebraic decay from transient exponential behaviour.
Outlook. Several extensions are natural: non–point initial data via the weight $W(\mathbf{q})$; two–point scalar statistics and fluxes; dependence on $\mathit{Sc}$ and weak inhomogeneity; stability of shells under weak forcing; and further rigor on the discrete–to–continuum limit (diffusion part) together with the momentum–space treatment of advection. On the validation side, targeted DNS that measure $U(\kappa)$ and, where feasible, the shell radii and jump law, would provide decisive tests of the scalar predictions derived here.
7 ACKNOWLEDGEMENTS
We gratefully acknowledge insightful discussions with Semon Rezchikov, Edoardo Spezzano, Katepalli Sreenivasan, and James Stone. The initial version of this work contained errors that affected the results. We are particularly grateful to an anonymous referee whose critical feedback significantly contributed to clarifying and correcting these issues. We extend an open invitation to this referee for further scientific dialogue via email or Zoom.
This research was partly supported by the Simons Foundation (award ID SFI-MPS-T-MPS-00010544) at the Institute for Advanced Study.
Declaration of Interests
The author reports no conflict of interest.
Declaration of AI use
I used an AI-assisted writing tool (ChatGPT, GPT-5) solely for stylistic editing and text formatting of this manuscript. The tool was employed to improve readability and flow, standardize terminology and section headings, and convert punctuation to LaTeX-safe forms. No scientific content, derivations, equations, symbols, figures, data analysis, or results were generated or altered by the tool. I manually reviewed and verified all edited text for accuracy and consistency with the submitted manuscript. Any errors remain my responsibility.