### v7

# Fluid dynamics, Statistical physics, Theoretical physics

Abstract
We derive an analytical solution for the density of a passive scalar in decaying homogeneous turbulence, considering the limit relevant to strong turbulence (high Reynolds number, fixed Schmidt number). The velocity statistics are governed by the Euler ensemble, which was previously obtained as a spontaneously stochastic exact solution of the loop equation derived from the Navier–Stokes equations in the strongly turbulent regime. The advection–diffusion problem for the scalar is also recast as a closed linear loop equation and solved by loop calculus.
For a localized initial condition, the solution consists of a sequence of expanding concentric shells: the radial scalar profile is quantized and piecewise parabolic, with gaps organized by Euler totients—an arithmetic structure distinct from conventional scaling predictions. This shell pattern is the unique solution within the Euler ensemble; finite diffusivity or sustained forcing smooths it. The result provides the underlying geometric structure for scalar transport in decaying strong turbulence and may be relevant in astrophysical or quantum-fluid regimes where dissipation is negligible. These findings may explain the so-called "ramp-cliff" structures observed in turbulent mixing half a century ago. While this shell structure is challenging to resolve directly in DNS, we show that its statistical signature is robustly captured by the volume-averaged scalar density, a quantity readily measurable in numerical simulations.
keywords:
Turbulence, Fractal, Transport Equation, Passive Scalar, Fixed Point, Velocity Circulation, Loop Equations
1 Preface: Microscopic theory of decaying turbulence
Turbulence —the complex, chaotic fluid motion observed everywhere, from water flowing from a faucet to the swirling atmospheres of stars —remains one of the oldest unsolved problems in classical physics. Despite extensive theoretical work and direct numerical simulations (DNS), a predictive, analytically tractable description has remained challenging. This paper presents an analytic solution for passive-scalar mixing in decaying turbulence. It builds on a recent solution for decaying turbulence derived analytically for the strongly turbulent regime, from first principles via loop calculus. In that formulation, the Navier-Stokes dynamics are recast in loop space, and a dual description emerges in which the turbulent regime is captured by the Euler ensemble, a spontaneously stochastic probability measure supported on solutions satisfying the Navier–Stokes loop equation. Discrete structures arise in momentum-loop space in this limit, with a regulator $N\to\infty$ and fixed “turbulent viscosity” $\tilde{\nu}=\nu N^{2}$.
Three conceptual shifts underlie the approach:
  1. 1.
    Loop-space reformulation. Rather than tracking the velocity field as a function of $(\mathbf{r},t)$, we study the evolution of circulation along closed loops in physical space. This yields a closed, linear equation for the loop functional-mathematically equivalent to Schrödinger (or complex diffusion) equation in loop space.
  2. 2.
    Spontaneous stochasticity in strong turbulence. The turbulent regime is described by a stochastic solution that covers a discrete ensemble of momentum loops, like the conservative Hamiltonian dynamics covers the energy surface. The solution applies in the limit $\nu\to 0$ with $\tilde{\nu}$ fixed, where a smooth loop is replaced by an $N$-gon and $N\to\infty$. The solution is realized as a random walk on a regular star polygon in momentum space.
  3. 3.
    Static and liquid loop equations. The classical (Eulerian) loop equation for a fixed loop has diffusion and advection terms. We also formulate the Lagrangian (“liquid”) loop equation for a loop advected by the flow; by Kelvin’s theorem, the advection term cancels and only the diffusion operator remains. We show that the same turbulent attractor solves both equations —the Euler ensemble.
  4. 4.
    Arithmetic structure. The resulting statistics exhibit an arithmetic organization (Euler totients and zeta sums) that will surface in the passive-scalar solution below.
DNS performed by Sreenivasan and collaborators already supports several predictions of the Euler ensemble solution in decaying turbulence. The predicted decay of turbulent kinetic energy, $\mathrm{TKE}(t)\sim t^{-5/4}$, agrees well with experiment and DNS, yielding a decay rate distinct from the Kolmogorov-Saffman prediction. Further agreement is seen in refined statistics such as the effective scaling exponent of the second-order velocity structure function as a function of the scaling variable $\log|\mathbf{r}|/\sqrt{t}$.
Here we extend the velocity solution to a second fundamental question: the transport of a passive scalar in decaying turbulence. Scalar mixing (temperature, concentration) is central in environmental, industrial, and astrophysical flows. Conventional treatments predict Gaussian spreading of a localized scalar, smoothed by molecular diffusion and eddy transport, but rely on surrogate velocity statistics or closure assumptions and encounter challenges at high Reynolds numbers.
The passive scalar does not feed back on the flow; it is advected by the velocity field and diffused molecularly. Mathematically, one seeks the scalar density as a functional of the velocity, then averaged over its statistics. Previous analytic results for scalar mixing often relied on surrogate velocity models (e.g., the Gaussian, $\delta$-correlated Kraichnan model [Falkovich et al.(2001)Falkovich, Gawȩdzki, and Vergassola]). The loop formulation presented here, however, yields a closed, linear equation for the scalar loop functional, which we solve directly at a fixed Schmidt number in the decaying regime.
We find that, in the Euler ensemble, a localized initial condition produces a hierarchy of expanding concentric shells: the radial scalar profile is quantized and piecewise parabolic, with gaps organized by Euler totients. This "shell" pattern represents the underlying geometric structure of scalar transport in ideal decaying turbulence; finite diffusivity or sustained forcing smooths it. The result is a parameter-free, analytic solution for passive-scalar statistics in the limit corresponding to strong turbulence. As we will show, while the sharp discontinuities are difficult to resolve, the statistical average of this structure within a volume is smooth. It provides a practical target for DNS verification.
Paper outline
Section 2 states the governing equations and the decaying-regime setting. Section 3 summarizes the loop calculus approach and the Euler ensemble solution for the velocity field, emphasizing the equivalence of static and liquid loops. Section 4 derives the passive-scalar loop equation and presents its analytic solution, including the shell structure. For a broader discussion relating these loop-space findings to classical turbulence paradigms, the reader is referred to Section 8 of our companion review [Migdal(2025a)]. Conclusions and implications are summarized in Section 5.
2 Introduction to the turbulent mixing
Turbulent mixing—the transport and diffusion of a passive scalar by a turbulent velocity field—is a fundamental process in many natural and technological systems. It governs the dispersion of pollutants, the distribution of heat and salinity in the oceans, and mixing in stars and industrial flows [Sreenivasan(2019), Dimotakis(2005), Warhaft(2000)]. The governing equations are the incompressible Navier–Stokes equations coupled to the scalar advection–diffusion equation:
$\frac{\partial\boldsymbol{v}}{\partial t}+(\boldsymbol{v}\cdot\nabla)\boldsymbol{v}=-\nabla p+\nu\nabla^{2}\boldsymbol{v},$ (1a)
$\nabla\cdot\boldsymbol{v}=0,$ (1b)
$\frac{\partial T}{\partial t}+(\boldsymbol{v}\cdot\nabla)T=D\nabla^{2}T,$ (1c)
where $\boldsymbol{v}(\boldsymbol{r},t)$ is the velocity field, $p$ is the pressure, $T(\boldsymbol{r},t)$ is the passive scalar, $\nu$ is the kinematic viscosity, and $D$ is the scalar diffusivity. The dimensionless Schmidt number is $\mathrm{Sc}=\nu/D$. Many theoretical, experimental, and numerical studies have addressed passive-scalar behavior, most often through global observables such as variance decay, spectra, or correlation functions (see, e.g., the review [Falkovich et al.(2001)Falkovich, Gawȩdzki, and Vergassola]). A key open question concerns the spatial structure that emerges from a localized initial condition in decaying turbulence. Traditional treatments suggest rapid Gaussian spreading under turbulent advection and molecular diffusion, but a closed-form solution for this regime has not been available.
We present an exact solution for the density of a passive scalar advected by a velocity field drawn from the Euler ensemble—a spontaneously stochastic probability measure supported on solutions of the unforced Navier–Stokes equations in the vanishing-viscosity limit [Migdal(2024)]. Using loop calculus, we solve the linear advection–diffusion equation without surrogate statistics or closure assumptions. The resulting scalar field exhibits a sequence of concentric spherical shells with quantized, piecewise-parabolic radial profiles and a smooth central value. This sharply ordered structure is a deterministic prediction of the ensemble and persists in the extreme turbulent limit; finite diffusivity smooths it. The effect may be relevant in regimes where dissipation is negligible, such as certain astrophysical or quantum-fluid settings.
3 Decaying turbulence and the Euler Ensemble
The underlying velocity statistics for the passive scalar mixing problem are taken from the recent exact solution of decaying turbulence derived via loop calculus [Migdal(2023), Migdal(2024), Migdal(2025c), Migdal(2025a)]. This approach reformulates the Navier-Stokes dynamics in terms of loop observables, leading to a solvable linear equation in loop space whose attractor, in the vanishing-viscosity limit, is the Euler ensemble [Migdal(2025a)]. We briefly summarize the key concepts relevant to the scalar problem.
3.1 Static vs. Liquid Loops and the Turbulent Attractor
Loop dynamics can be described in either an Eulerian frame (static loop $C$) or a Lagrangian frame (liquid loop advected with the flow). The circulation
$\Gamma_{C}[v]=\oint_{c}d\boldsymbol{r}\cdot\boldsymbol{v}(\boldsymbol{r},t)$
evolves differently in each frame. In the Eulerian frame
$\partial_{t}\Gamma=\oint_{C}d\boldsymbol{r}\cdot\left(-\nu\boldsymbol{\nabla}\times\boldsymbol{\omega}+\boldsymbol{v}\times\boldsymbol{\omega}\right);$ (2)
$\omega=\boldsymbol{\nabla}\times v;$ (3)
In the Lagrangian frame, Kelvin’s theorem ensures that the advection term $\boldsymbol{v}\times\boldsymbol{\omega}$ is canceled by the loop’s motion, leaving only the viscous diffusion term:
$\partial_{t}\boldsymbol{C}(\theta,t)=\boldsymbol{v}\left(\boldsymbol{C}(\theta,t),t\right);$ (4)
$\partial_{t}\Gamma_{C}[v]=-\nu\oint_{C}d\boldsymbol{r}\cdot\boldsymbol{\nabla}\times\boldsymbol{\omega};$ (5)
While distinct, both the static and liquid loop equations derived from Navier-Stokes are solved by the same universal turbulent attractor state in the decaying, vanishing-viscosity limit [Migdal(2023), Migdal(2025a)]. This attractor, the Euler ensemble $\mathcal{E}$, corresponds to a uniform measure over a discrete set of momentum loop configurations. Crucially, the Euler ensemble solution satisfies the liquid loop equation (where advection is absent by Kelvin’s theorem). It has been shown explicitly that the advection term $\mathbf{v}\times\boldsymbol{\omega}$ integrates to zero around the loop for this specific solution, thus satisfying the static loop equation as well (see Subsection 3(h) and Appendix B of [Migdal(2025a)]). This equivalence allows us to use the simpler liquid-loop framework, knowing that its solution applies to the physical Eulerian frame.
3.2 Loop Calculus Framework and the Euler Ensemble Solution
The mathematical framework underlying this solution is the loop space calculus, which uses functional derivatives with respect to the loop’s velocity profile $\mathbf{C}^{\prime}(\theta)$ to represent differential operators and field strengths acting on loop functionals [Makeenko and Migdal(1979), Migdal(1983), Migdal(1998), Migdal(2025b), Migdal(2025a)]. The details of this calculus are presented in Appendix A of [Migdal(2025a)]. This approach avoids the singularities of earlier methods and allows for the derivation of closed, linear evolution equations for the loop functional $\Psi[C,t]=\langle\exp\left(\mathrm{i}\Gamma_{C}[v]/\nu\right)\rangle_{v}$.
The liquid loop equation takes the form of a diffusion equation for $\Psi[C]$:
$\partial_{t}\Psi[C,t]=\nu\mathcal{L}_{C}\Psi[C,t]$ (6)
with the diffusion operator
$\mathcal{L}_{C}=\oint d\theta C^{\prime}_{\nu}(\theta)\hat{L}_{\nu}(\theta);$ (7)
$\hat{L}_{\nu}(\theta)=T^{\alpha\beta\gamma}_{\nu}\frac{\delta^{3}}{\delta C^{\prime}_{\alpha}(\theta-0)\delta C^{\prime}_{\beta}(\theta)\delta C^{\prime}_{\gamma}(\theta+0)};$ (8)
$T^{\alpha\beta\gamma}_{\nu}=\delta_{\alpha\beta}\delta_{\gamma\nu}+\delta_{\gamma\beta}\delta_{\alpha\nu}-2\delta_{\alpha\gamma}\delta_{\beta\nu};$ (9)
All three arguments of the third functional derivative tend to $\theta$ in the specified order.
The result of this functional differentiation depends solely on the point order and is purely kinematical and unrelated to any potential singularities in the velocity or vorticity fields. These discontinuities are present even for a pure rotation where $\boldsymbol{\omega}$ does not depend on coordinates. The origin of discontinuity is the existence of vorticity: the components of the second functional derivative $\frac{\delta}{\delta C^{\prime}_{\alpha}(\theta-0)}\frac{\delta}{\delta C^{\prime}_{\beta}(\theta+0)}\Gamma_{C}[v]$ create different components of the antisymmetric tensor $\omega_{\alpha\beta}=\partial_{\alpha}v_{\beta}-\partial_{\beta}v_{\alpha}$. This asymmetry leads to a nontrivial limit as two points tend to the same $\theta$, so the result changes sign under a different ordering of points, $\theta\pm$. This is how the 3D rotation is reflected in one-dimensional loop kinematics: through discontinuities in higher-order functional derivatives.
3.3 The Momentum loop equation, and the meaning of discontinuities
This equation is solved via a functional Fourier transform to "momentum loop space," leading to a nonlinear equation for the momentum loop $P(\theta,t)$.
$\Psi[C,t]=\left\langle\exp\left(\mathrm{i}\int_{0}^{2\pi}d\theta\boldsymbol{C}^{\prime}_{\mu}(\theta)\cdot\boldsymbol{P}(\theta,t)\right)\right\rangle_{P(t)};$ (10)
$\frac{1}{\nu}\partial_{t}\boldsymbol{P}(\theta)=(\Delta\boldsymbol{P}(\theta)\cdot\boldsymbol{P}(\theta))\Delta\boldsymbol{P}(\theta)-(\Delta\boldsymbol{P}(\theta))^{2}\boldsymbol{P}(\theta);$ (11)
$\boldsymbol{P}(\theta)\equiv\frac{\boldsymbol{P}(\theta+)+\boldsymbol{P}(\theta-)}{2};$ (12)
$\Delta\boldsymbol{P}(\theta)\equiv\boldsymbol{P}(\theta+)-\boldsymbol{P}(\theta-);$ (13)
The brackets $\left\langle\right\rangle_{P(t)}$ correspond to the averaging over an ensemble of solutions of the time evolution of $P_{\mu}(.,t)$ described by the above equation.
The discontinuity of the momentum loop $\Delta\boldsymbol{P}(\theta)$ reflects the ordering of derivatives in the diffusion operator (7). As mentioned above, the local vorticity at the loop is represented by the cross product of left and right derivatives $\frac{\delta}{\delta\boldsymbol{C}^{\prime}(\theta-)}\times\frac{\delta}{\delta\boldsymbol{C}^{\prime}(\theta+)}$. In Fourier loop space, this becomes the cross product $\boldsymbol{P}(\theta-)\times\boldsymbol{P}(\theta+)=\boldsymbol{P}(\theta)\times\Delta\boldsymbol{P}(\theta)$. Thus, the flow with vorticity can only be described in Fourier loop space by momentum loops with discontinuities at every angle. These functions can only be defined as distributions, as demonstrated in the original papers [Migdal(2023), Migdal(2024)] using the example of a uniform random rotation. In that example, the discontinuities were manifested in the slow convergence of a Fourier series of a periodic function $\boldsymbol{P}(\theta)$. The need to work with such unusual functions is a major obstacle to the acceptance of momentum loop dynamics within the turbulence community, which is accustomed to continuous functions.
In the Euler ensemble below, these discontinuities have a different, number-theoretical origin and lead to a nontrivial vorticity distribution.
3.4 The Euler ensemble as solution of the Momentum loop equation
In the decaying turbulence limit ($\nu\to 0$ with $\tilde{\nu}=\nu N^{2}$ fixed, where $N$ is a loop discretization parameter tending to infinity), the attractor solution for the momentum loop corresponds to the Euler ensemble $\mathcal{E}$, as derived in Sections 3-4 of [Migdal(2025a)] (see also [Migdal(2023), Migdal(2024)]). The time dependence is a trivial consequence of power counting on both sides of the (11), leading to algebraic equations for the emerging universal vector function $\boldsymbol{F}(\theta)$
$\boldsymbol{P}(\theta)=\frac{\boldsymbol{F}(\theta)}{\sqrt{2\nu\,(t+t_{0})}};$ (14)
$\left(\Delta\boldsymbol{F}^{2}-1\right)\boldsymbol{F}=(\Delta\boldsymbol{F}\cdot\boldsymbol{F})\Delta\boldsymbol{F};$ (15)
$\boldsymbol{F}\equiv\frac{\boldsymbol{F}(\theta+)+\boldsymbol{F}(\theta-)}{2};$ (16)
$\Delta\boldsymbol{F}\equiv\boldsymbol{F}(\theta+)-\boldsymbol{F}(\theta-);$ (17)
The manifold $\mathcal{E}$ of solutions of these algebraic equations for $\mathbf{F}(\theta)$ is the set $\mathcal{E}$ of random walks on regular star polygons $\{q/p\}$, which we named the Euler ensemble.
The detailed derivation and properties are found in [Migdal(2024)] and Section 4 of [Migdal(2025a)]. The crucial point is that $\mathcal{E}$ provides a specific, analytically defined statistical ensemble for the velocity field in the turbulent limit. The discontinuities at every angle are derived as a limit of the discrete values $\boldsymbol{F}_{k}=\boldsymbol{F}(\theta_{k}),\theta_{k}=2\pi k/N$. Here we reproduce this solution for the readers’ convenience in the current notation.
The Euler Ensemble: A Universal Turbulent Attractor
In this solution, $\boldsymbol{F}(\theta)$ is a universal fractal curve constructed as the continuum limit ($N\to\infty$) of a random walk on a regular star polygon $\{q/p\}$. Its vertices are given by:
$\boldsymbol{F}\left(\frac{2\pi k}{N}\right)=\hat{\Omega}\cdot\{R\cos(\alpha_{k}),R\sin(\alpha_{k}),0\};$ (18)
$R=\frac{1}{2\sin(\beta/2)};$ (19)
where the rotation matrix $\hat{\Omega}\in SO(3)$, the angle step $\beta=2\pi p/q$ with $p/q\in\mathbb{Q}$, and the cumulative angle $\alpha_{k}=\beta\sum_{l=1}^{k}\sigma_{l}$ with $\sigma_{l}=\pm 1$ define the random walk.
The recurrent equations (15) are satisfied in an obvious geometric way: with this radius $R$ the sides have unit length:
$R\left|e^{\mathrm{i}\alpha_{k+1}}-e^{\mathrm{i}\alpha_{k}}\right|=R\left|e^{\mathrm{i}\beta\sigma_{k}}-1\right|=R\left|2\mathrm{i}\sin(\beta/2)\right|=1.$
Due to this property, $|\Delta\boldsymbol{F}|=1,\boldsymbol{F}\cdot\Delta\boldsymbol{F}=0$, which nullifies both sides of the (15). The Euler ensemble is not a single configuration but a degenerate manifold of solutions, a "fixed stochastic trajectory." The statistical state of turbulence is a distribution over all choices of $(\hat{\Omega},p,q,\{\sigma_{k}\})$.
The periodicity condition $\alpha_{N}-\alpha_{0}=\beta\sum\sigma_{k}=2\pi r$ imposes a number-theoretic constraint on the geometry of the walk: for a walk to close after traversing N steps, the rotation angle $\beta$ must be a rational multiple of $2\pi$, i.e., $\beta=2\pi p/q$. The requirement that the polygon $\{q/p\}$ be irreducible (i.e., not a simpler polygon traced multiple times) restricts $p$ and $q$ to be coprime, meaning the number of unique star polygons for a given denominator $q$ is precisely $\varphi(q)$, the Euler totient function.
The statistics of turbulence are thus described by a uniform distribution over all such polygons and all possible random walks on them. This discrete, number-theoretic structure is the hallmark of spontaneous quantization. This discrete distribution is a crucial point– the origin of the link between turbulence and number theory. We are limited to the manifold $\mathbb{Q}$ of rational numbers, which eventually leads to all the "quantum" effects in the solution of decaying turbulence, including the discrete shell structure of passive scalar density we derive below.
The Euler ensemble is remarkably universal. As we show in the Appendix of [Migdal(2025a)], it also describes the solution of the loop equation for the static loop (not moving with the flow). Explicit computation shows that the advection term $\vec{v}\times\vec{\omega}$ present in the static loop equation, when integrated over the loop using the solution for vorticity and (by inverting the curl operator) for velocity for the Euler ensemble, becomes a total derivative of a periodic function, so that the integral over $\theta$ yields zero. This nontrivial cancellation was not built into the solution for the liquid loop but emerged as a free bonus, suggesting a deeper reason for this solution – not yet completely understood.
3.5 Connection to Observables and Validation
This loop-based solution allows the calculation of physical observables. For instance, vorticity correlations can be obtained via functional derivatives of $\Psi[C,t]$ evaluated on specific loop shapes like the hairpin loop. The Euler ensemble yields explicit predictions for these correlations, as well as for the turbulent kinetic energy decay ($E(t)\sim t^{-5/4}$) and the integral scale growth ($L(t)\sim t^{1/2}$), detailed in Section 4.1 of [Migdal(2025a)] (see also [Migdal(2024)]). These parameter-free predictions agree well with experimental data and high-resolution DNS, particularly in the decaying regime, where classical models like the Kolmogorov-Saffman model struggle [Comte-Bellot and Corrsin(1966), Panickacheril John et al.(2022a)Panickacheril John, Donzis, and Sreenivasan, Panickacheril John et al.(2022b)Panickacheril John, Donzis, and Sreenivasan]. This validation supports the use of the Euler ensemble as the correct statistical description of the velocity field for the subsequent passive scalar analysis.
In what follows, we take this validated Euler ensemble solution for the velocity field as given and apply its statistics to solve the passive-scalar transport problem.
4 Passive scalar dynamics in loop space
In the conventional approach [Falkovich et al.(2001)Falkovich, Gawȩdzki, and Vergassola], the advection–diffusion equation is coupled to an unknown velocity field; progress typically comes from Gaussian surrogates or ad hoc closures of the moment hierarchy. In the loop formulation, the problem reduces to a closed linear equation for the loop functional, which we solve exactly—no surrogates, no closures.
4.1 circulation and temperature joint distribution
Consider the passive scalar (temperature) evaluated at a loop point $\boldsymbol{r}_{1}=\boldsymbol{C}(\theta_{1})$. The advection–diffusion equation can be written in terms of loop operators as
$\partial_{t}T(\boldsymbol{C}(\theta_{1}),t)=\left(-\boldsymbol{v}(\boldsymbol{C}(\theta_{1}))\cdot\partial_{\boldsymbol{C}(\theta_{1})}+D\left(\partial_{\boldsymbol{C}(\theta_{1})}\right)^{2}\right)T(\boldsymbol{C}(\theta_{1}),t)$ (20)
Note that the gradient $\partial_{\boldsymbol{C}(\theta_{1})}$ represents the singular limit in which only a single point on the loop is shifted
$\partial_{\boldsymbol{C}(\theta)}=\frac{\delta}{\delta\boldsymbol{C}^{\prime}(\theta-0)}-\frac{\delta}{\delta\boldsymbol{C}^{\prime}(\theta+0)};$ (21)
The last formula for the gradient was proven in loop space calculus (Appendix A of [Migdal(2025a)]) using a conventional functional derivative
$\partial_{\boldsymbol{C}(\theta)}=\int_{\theta-0}^{\theta+0}d\theta^{\prime}\frac{\delta}{\delta\boldsymbol{C}(\theta^{\prime})}=-\int_{\theta-0}^{\theta+0}d\theta^{\prime}\partial_{\theta^{\prime}}\frac{\delta}{\delta\boldsymbol{C}^{\prime}(\theta^{\prime})}=\frac{\delta}{\delta\boldsymbol{C}^{\prime}(\theta-0)}-\frac{\delta}{\delta\boldsymbol{C}^{\prime}(\theta+0)};$ (22)
In momentum loop space, this discontinuity of the functional derivative translates to a discontinuity of the momentum loop.
Because the scalar does not feed back on the flow, the Euler ensemble continues to describe the velocity statistics. Diffusion affects only the scalar statistics, so we must solve the advection–diffusion problem in loop space given by (20), and then average the resulting solution over the distribution of the velocity field that follows from the turbulent solution of the Navier-Stokes equation (1). This task seems impossible within the direct approach to these equations, but with the loop space approach, these obstacles can be bypassed.
Let us consider the mixed average
$T_{C}(\theta_{1},t)=\left\langle T\left(\boldsymbol{C}(\theta_{1}),t\right)\exp\left(\frac{\mathrm{i}\Gamma_{C}[v]}{\nu}\right)\right\rangle_{v}$ (23)
which correlates the temperature at $\boldsymbol{C}(\theta_{1})$ with the circulation around $C$. Unlike the mean temperature $\left\langle T(\vec{r},t)\right\rangle_{v}$, which is a function of a single point in space (plus time), this $T_{C}(\theta_{1},t)$ is a functional of the loop $C$. This step is the first in our ascent to the loop space from the physical $3+1$-dimensional space.
Differentiating in time produces contributions from both (20) and the Navier-Stokes dynamics:
$\partial_{t}T_{C}(\theta_{1},t)=\left\langle\exp\left(\frac{\mathrm{i}\Gamma_{C}[v]}{\nu}\right)\left(-\boldsymbol{v}(\boldsymbol{C}(\theta_{1}))\cdot\partial_{\boldsymbol{C}(\theta_{1})}+D\partial_{\boldsymbol{C}(\theta_{1})}^{2}\right)T\left(\boldsymbol{C}(\theta_{1}),t\right)\right\rangle_{v}$
$+\nu\left\langle T\left(\boldsymbol{C}(\theta_{1}),t\right)\mathcal{L}_{C}\exp\left(\frac{\mathrm{i}\Gamma_{C}[v]}{\nu}\right)\right\rangle_{v}$ (24)
We are going to take advantage of the fact established in Appendix A of [Migdal(2025a)] that the advection term cancels in the Euler ensemble solution for the Navier-Stokes loop equation, so we were left only with the diffusion term in loop operator $\mathcal{L}_{C}$, (6).
4.2 The loop equation for the passive scalar
The velocity and vorticity fields are represented in loop space as operators
$\hat{\omega}(\theta)=-\mathrm{i}\nu\frac{\delta}{\delta\boldsymbol{\sigma}(\theta)}=-\mathrm{i}\nu\frac{\delta}{\delta\boldsymbol{C}^{\prime}(\theta-)}\times\frac{\delta}{\delta\boldsymbol{C}^{\prime}(\theta+)};$ (25)
$\hat{v}(\theta_{1})=\frac{-1}{\partial_{\boldsymbol{C}(\theta_{1})}^{2}}\partial_{\boldsymbol{C}(\theta_{1})}\times\hat{\omega}(\theta);$ (26)
In order to rewrite (4.1) as a loop equation for $T_{C}$, we need to extend the operators on all the factors inside averaging $\left\langle\right\rangle_{v}$ over Navier-Stokes dynamics. Taking the loop operators outside the average $\left\langle\cdot\right\rangle_{v}$ would, in principle, generate cross terms when $\partial_{\boldsymbol{C}(\theta_{1})}$ acts on $\exp\left(\frac{\mathrm{i}\Gamma_{C}[v]}{\nu}\right)$ and when $\mathcal{L}_{C}$ acts on $T(\boldsymbol{C}(\theta_{1}),t)$. All such cross terms vanish by the loop calculus identities proven in Appendix A, sections (h) and (i) of [Migdal(2025a)]. Thus, we get a linear diffusion-advection equation in loop space
$\partial_{t}T_{C}(\theta_{1},t)=\left(-\boldsymbol{v}(\boldsymbol{C}(\theta_{1}))\cdot\partial_{\boldsymbol{C}(\theta_{1})}+D\partial_{\boldsymbol{C}(\theta_{1})}^{2}+\nu\mathcal{L}_{C}\right)T_{C}(\theta_{1},t)\$(27)
4.3 Solution of the passive scalar loop equation
The solution for the above equation has the form of an average over the Euler ensemble
$T_{C}(\theta_{1},t)=\left\langle\tau(\boldsymbol{C}(\theta_{1}),t)\exp\left(\frac{\mathrm{i}\oint d\theta\boldsymbol{C}^{\prime}(\theta)\cdot\boldsymbol{F}(\theta)}{\sqrt{2\nu(t+t_{0})}}\right)\right\rangle_{\mathcal{E}}$ (28)
where $\tau$ is determined by
$\frac{\partial\tau(\boldsymbol{C}(\theta_{1}),t)}{\partial t}\exp\left(\frac{\mathrm{i}\oint d\theta\boldsymbol{C}^{\prime}(\theta)\cdot\boldsymbol{F}(\theta)}{\sqrt{2\nu(t+t_{0})}}\right)=$
$\left(-\hat{v}(\theta_{1})\cdot\partial_{\boldsymbol{C}(\theta_{1})}+D\partial_{\boldsymbol{C}(\theta_{1})}^{2}\right)\left(\tau(\boldsymbol{C}(\theta_{1}),t)\exp\left(\frac{\mathrm{i}\oint d\theta\boldsymbol{C}^{\prime}(\theta)\cdot\boldsymbol{F}(\theta)}{\sqrt{2\nu(t+t_{0})}}\right)\right);$ (29)
and the remaining term $\nu\mathcal{L}_{C}$ in (27) cancel by the momentum loop equation (11) for $\boldsymbol{P}=\boldsymbol{F}/\sqrt{2\nu(t+t_{0})}$.
With extra terms coming from applying the operator $\partial_{\boldsymbol{C}(\theta_{1})}$ to the exponential, we obtain a linear evolution for $\tau$,
$\frac{\partial\tau(\boldsymbol{C}(\theta_{1}),t)}{\partial t}=\hat{Q}(\theta_{1},t)\tau(\boldsymbol{C}(\theta_{1}),t);$ (30)
$\hat{Q}(\theta,t)=\left(\frac{-2\mathrm{i}D\Delta\boldsymbol{F}(\theta)}{\sqrt{2\nu(t+t_{0})}}-\tilde{v}(\theta,t)\right)\cdot\partial_{\boldsymbol{C}(\theta)}+D\partial_{\boldsymbol{C}(\theta)}^{2}-D\frac{(\Delta\boldsymbol{F}(\theta))^{2}}{2\nu(t+t_{0})};$ (31)
$\tilde{v}(\theta,t)=\left.\hat{v}(\theta_{1},t)\right|_{\mathcal{E}}$ (32)
Here $\left.\hat{v}(\theta_{1},t)\right|_{\mathcal{E}}$ is the velocity operator in momentum loop space, with momentum loop according to the Euler ensemble
$\left.\hat{v}(\theta_{1})\right|_{\mathcal{E}}=\frac{1}{\Delta\boldsymbol{F}^{2}(\theta_{1})}\Delta\boldsymbol{F}(\theta_{1})\times(\boldsymbol{F}(\theta_{1})\times\Delta\boldsymbol{F}(\theta_{1}))\frac{\nu}{\sqrt{2\nu(t+t_{0})}}=\frac{\nu\boldsymbol{F}(\theta_{1})}{\sqrt{2\nu(t+t_{0})}};$ (33)
4.4 The discretization of the solution
The strict mathematical meaning of these discontinuities can be provided by replacing the loop $C$ with a polygon, corresponding to the equidistant angles $\theta_{k}=2\pi k/N$ on a unit circle. This discretization was used in [Elia and Lellis(2025)] to justify the loop calculus and prove that the Euler ensemble solves it at finite $N$ (theorem 10.6 in their paper). This regulator must tend to infinity in final observables, as is done in quantum field theory.
Then, with $\boldsymbol{F}_{k}\equiv\boldsymbol{F}(\theta_{k})$
$\Delta\boldsymbol{F}(\theta_{k})=\boldsymbol{F}_{k}-\boldsymbol{F}_{k-1};$ (34)
With $|\Delta\boldsymbol{F}(\theta_{k})|^{2}=1$ one finds
$\tilde{v}(\theta_{k},t)=\frac{\nu\boldsymbol{F}_{<k>}}{\sqrt{2\nu(t+t_{0})}};$ (35)
$\boldsymbol{F}_{<k>}=\frac{\boldsymbol{F}_{k}+\boldsymbol{F}_{k-1}}{2};$ (36)
This simple differential equation is solvable in the Fourier representation
$\tau(\boldsymbol{r},t)=\int\frac{d^{3}q}{(2\pi)^{3}}\exp\left(\mathrm{i}\boldsymbol{q}\cdot\boldsymbol{r}\right)\eta(\boldsymbol{q},t);$ (37)
Substituting this Anzatz into (30) we get linear ODE for $\eta(\boldsymbol{q},t)$
$\frac{d\eta(\boldsymbol{q},t)}{dt}=\left(\boldsymbol{q}\cdot\frac{2D\Delta\boldsymbol{F}_{k}-\mathrm{i}\nu\boldsymbol{F}_{<k>}}{\sqrt{2\nu(t+t_{0})}}-D\boldsymbol{q}^{2}-\frac{D}{2\nu(t+t_{0})}\right)\eta(\boldsymbol{q},t)$ (38)
which we readily solve:
$\eta(\boldsymbol{q},t)=W(\boldsymbol{q})\left(\frac{t+t_{0}}{t_{0}}\right)^{-\frac{1}{2\mathit{Sc}}}\exp\left(-\frac{\boldsymbol{q}^{2}\nu t}{\mathit{Sc}}-\mathrm{i}\left(\sqrt{2\nu(t+t_{0})}-\sqrt{2\nu t_{0}}\right)\boldsymbol{q}\cdot\boldsymbol{G}_{k}\right);$ (39)
$\boldsymbol{G}_{k}=\frac{\boldsymbol{F}_{k}+\boldsymbol{F}_{k-1}}{2}+2\mathrm{i}\frac{\boldsymbol{F}_{k}-\boldsymbol{F}_{k-1}}{\mathit{Sc}}$ (40)
Here $\mathit{Sc}=\nu/D$ is the Schmidt number, the sole dimensionless parameter for the scalar.
The weight $W(\boldsymbol{q})$ reflects the initial scalar distribution:
$W(\boldsymbol{q})=\eta(\boldsymbol{q},0)=\int d^{3}r\exp\left(-\mathrm{i}\boldsymbol{q}\cdot\boldsymbol{r}\right)T(\boldsymbol{r},0)$ (41)
The Fourier quantity $\eta(\boldsymbol{q},t)$ is complex because the temperature is averaged with the phase factor $\exp\left(\mathrm{i}\Gamma_{C}[v]/\nu\right)$.
The mean temperature is, of course, real and positive. To extract it, contract the loop to a point: set $\boldsymbol{r}=\boldsymbol{C}(\theta_{1})$ and $\boldsymbol{C}^{\prime}(\theta)=0$. Then $\Gamma_{C}[v]=0$, and the original average (23) reduces to the scalar mean at a point. On the right of (28), setting $\boldsymbol{C}^{\prime}=0$ eliminates the momentum-loop factor. Thus
$\left\langle T(\boldsymbol{r},t)\right\rangle_{v}=\left\langle\tau(\boldsymbol{C}(\theta_{1}),t)\right\rangle_{\mathcal{E}}=\left(\frac{t+t_{0}}{t_{0}}\right)^{-\frac{1}{2\mathit{Sc}}}$
$\int\frac{d^{3}q}{(2\pi)^{3}}W(\boldsymbol{q})\exp\left(\mathrm{i}\boldsymbol{q}\cdot\boldsymbol{r}\right)\left\langle\exp\left(-\frac{\boldsymbol{q}^{2}\nu t}{\mathit{Sc}}-\mathrm{i}\left(\sqrt{2\nu(t+t_{0})}-\sqrt{2\nu t_{0}}\right)\boldsymbol{q}\cdot\boldsymbol{G}_{k}\right)\right\rangle_{\mathcal{E}};$ (42)
4.5 The turbulent limit. Harmony of primes behind chaos
As shown in [Migdal(2023), Migdal(2024)], the turbulent limit combines infinitesimal viscosity with large polygon radius:
$\nu\to\frac{\tilde{\nu}}{N^{2}}=O(1/N^{2});$ (43)
$R\to\rho N=O(N);$ (44)
$\boldsymbol{F}(\theta)\to N\rho\left\{\cos\alpha,\sin\alpha,0\right\};$ (45)
$|\Delta\boldsymbol{F}(\theta)|=1;$ (46)
In this limit, the vector $\sqrt{\nu}\boldsymbol{G}_{k}$ tends to a finite real vector,
[…]
The distribution of $\rho=\big(2N\sin(\pi p/q)\big)^{-1}$ for coprime $p,q$ with $0<p<q<N\to\infty$ was found in [Migdal(2024)] by some advanced methods of number theory. This is a discontinuous, piecewise power-like distribution:
$f(\rho)=\left(1-\frac{\pi^{2}}{21600\zeta(5)}\right)\delta(\rho)+\frac{2\pi^{3}}{3}\rho^{4}\Phi\left(\left\lfloor\frac{1}{2\pi\rho}\right\rfloor\right);$ (48)
with $\Phi$ the totient summatory function:
$\Phi(q)=\sum_{n=1}^{q}\varphi(n);$ (49)
$\varphi(m)=m\prod_{p|m}\left(1-\frac{1}{p}\right)$ (50)
For a point-source initial condition,
$T(\boldsymbol{r},0)=T_{0}\delta(\boldsymbol{r})$ (51)
we find from (37)
$\left\langle T(\boldsymbol{r},t)\right\rangle_{v}=T_{0}(1+t/t_{0})^{-\frac{1}{2\mathit{Sc}}}G(r,t);$ (52)
$G(r,t)=\left\langle\delta\left(\boldsymbol{r}-\rho\left(\sqrt{2\tilde{\nu}(t+t_{0})}-\sqrt{2\tilde{\nu}t_{0}}\right)\hat{\Omega}\cdot\left\{\cos\alpha_{k},\sin\alpha_{k},0\right\}\right)\right\rangle_{\rho,\hat{\Omega}}$ (53)
The one-dimensional delta in (48) produces a three-dimensional delta in $G$; the remaining term yields a finite structure. Averaging $G(r,t)$ over $\hat{\Omega}\in\mathbb{O}(3)$ removes the dependence upon $\alpha_{k}$ leading to a simple answer
$G(r,t)=\left(1-\frac{\pi^{2}}{21600\zeta(5)}\right)\delta(\boldsymbol{r})+\frac{\pi^{2}}{6r^{2}}\int d\rho\rho^{4}\delta\left(r-2\pi\rho L(t)\right)\Phi\left(\left\lfloor\frac{1}{2\pi\rho}\right\rfloor\right);$ (54)
$L(t)=\frac{\sqrt{2\tilde{\nu}(t+t_{0})}-\sqrt{2\tilde{\nu}t_{0}}}{2\pi};$ (55)
and after integrating over $\rho$,
$\left\langle T(\boldsymbol{r},t)\right\rangle_{v}=(1+t/t_{0})^{-\frac{1}{2\mathit{Sc}}}\left[\left(1-\frac{\pi^{2}}{21600\zeta(5)}\right)\delta(\boldsymbol{r})+\frac{r^{2}\Phi\left(\left\lfloor\frac{L(t)}{r}\right\rfloor\right)}{768\pi^{5}L(t)^{5}}\right];$ (56)
This is a decaying prefactor multiplying a rescaled delta function plus a second, decaying contribution with a discontinuous dependence on the scaling variable $r/L(t)$ (Fig. 1). The support lies at $r\leq L(t)$, so the shells contract towards the origin as $t\downarrow 0$.
Thus, a point source produces infinitely many concentric spherical shells, $\frac{L(t)}{n}<r<\frac{L(t)}{n-1}$, accumulating at the origin. The profile is parabolic within each shell, with jumps $\propto\varphi(n)/n^{2}$ at successive boundaries. Since these jumps decrease with $n$, the temperature approaches a finite limit at the origin:
$\lim_{n\to\infty}\frac{\Phi\left(n\right)}{n^{2}}=\frac{3}{\pi^{2}}$ (57)
The discontinuities reflect the absence of a continuum limit for Euler’s totient; e.g.,
$\varphi(7919)=7918;$ (58)
$\varphi(7920)=1920;$ (59)
although statistically, for large $n$,
$\varphi(n)\sim\textit{ const }{}\frac{n}{\log\log n}\quad\text{for almost all }n$ (60)
Observing such a shell structure directly in DNS would require very large grids; in physical flows, it may be smoothed by weak dissipation or forcing. Predicted structure of discrete shells with sharp interfaces might explain the ’ramp-cliff’ patterns observed experimentally in passive scalar fields in turbulent flows[Sreenivasan(2019)] (see also Fig. 5 in [Migdal(2025a)]), which feature similar asymmetric profiles.
Figure 1: Log–log plot of the universal function $\xi^{2}\Phi\left(\left\lfloor\frac{1}{\xi}\right\rfloor\right)$ where $\xi=\frac{2\pi r}{\left(\sqrt{2\tilde{\nu}(t+t_{0})}-\sqrt{2\tilde{\nu}t_{0}}\right)}$.
For spectral properties, take $W(\boldsymbol{q}^{\prime})=\delta(\boldsymbol{q}-\boldsymbol{q}^{\prime})$ (a single Fourier mode):
$\left\langle\tilde{T}(\boldsymbol{q},t)\right\rangle_{v}=T_{0}\left(1+\frac{t}{t_{0}}\right)^{-\frac{1}{2\mathit{Sc}}}U\left(|\boldsymbol{q}|\left(\sqrt{2\tilde{\nu}(t+t_{0})}-\sqrt{2\tilde{\nu}t_{0}}\right)\right);$ (61)
$U(\kappa)=\left\langle\exp\left(\mathrm{i}\rho\boldsymbol{\kappa}\cdot\hat{\Omega}\cdot\left\{\cos\alpha_{k},\sin\alpha_{k},0\right\}\right)\right\rangle_{\rho,\hat{\Omega}}=\left\langle\frac{\sin(\rho\kappa)}{\rho\kappa}\right\rangle_{\rho}$ (62)
Using (48) and the Taylor expansion of $\sin(\rho\kappa)/(\rho\kappa)$ yields an absolutely convergent series
$U(\kappa)=1+\sum_{l=1}^{\infty}\frac{(l+1)(2\pi)^{-2l-3}\left(-\kappa^{2}\right)^{l}\zeta(2(l+2))}{3(2l+5)\zeta(2l+5)\Gamma(2l+3)}$ (63)
We have evaluated this series numerically in Mathematica ® with high precision: $U(\kappa)$ starts at $1$, decreases slightly, and then oscillates about $0.99986$ (Fig. 2).
Figure 2: The plot of universal function $U(\kappa)$ in (63) describing the amplitude of passive scalar plane wave with wavevector $\boldsymbol{q}$ at the time $t$ as a function of scaling variable $\kappa=|\boldsymbol{q}|\left(\sqrt{2\tilde{\nu}(t+t_{0})}-\sqrt{2\tilde{\nu}t_{0}}\right)$
4.6 How did the loop space solve an impossible problem?
This approach provides an analytic solution to the passive scalar problem, a task that appears intractable when approached directly. A direct calculation would require averaging the scalar field $T(\mathbf{r},t)$, which is an unknown functional of the velocity $\mathbf{v}$, over the unknown statistical distribution of $\mathbf{v}$ itself. How, then, is this seemingly insurmountable difficulty overcome?
We circumvented the direct problem by moving to loop space. Instead of focusing on the velocity field directly, we leverage the known diffusion laws governing the evolution of arbitrary liquid loops advected by the flow. We consider the scalar value $T$ at a specific point on such a loop, passively moving with it.
The crucial step is to define a joint loop functional that incorporates both the scalar value at a point and the circulation around the entire loop, $\left\langle T(\boldsymbol{C}(\theta_{1}),t)\exp\left(\frac{\mathrm{i}\Gamma_{C}[v]}{\nu}\right)\right\rangle_{v}$. When the governing equations for both the scalar (advection-diffusion) and the velocity (Navier-Stokes, expressed via circulation dynamics) are transformed into loop space operators acting on this joint functional, a remarkable simplification occurs: the resulting evolution equation becomes linear.
Within this linear functional equation, the statistical properties of the velocity field, as captured by the Euler ensemble solution $\Psi[C,t]$ for the circulation functional, effectively become part of the known linear operator. Solving this equation, via functional Fourier transforms (leading to momentum loop space), yields the evolution of the joint functional. The desired average scalar value $\left\langle T(\mathbf{r},t)\right\rangle_{v}$ is then recovered by evaluating this solution in the limit of a contracted loop (equivalent to setting $\boldsymbol{C}^{\prime}(\theta)=0$). Note that in this limit, the phase factor $\exp\left(\mathrm{i}\frac{\Gamma_{C}}{\nu}\right)$ tends to $1$, and so does the factor $\exp\left(\mathrm{i}\oint d\theta\boldsymbol{C}^{\prime}(\theta)\cdot\boldsymbol{P}(\theta,t)\right)$. The role of these factors was to participate in time differentiation and loop differentiation while the loop was still finite. In the end, the factors disappeared, but their contribution to the loop equation remained. This juggling of the loop, which after differentiation shrinks to zero circulation, is the standard technique used in loop calculus to obtain ordinary correlation functions [Migdal(2024)].
Essentially, the knowledge of the velocity distribution and the functional dependence of $T$ on $\mathbf{v}$ is implicitly encoded within the loop diffusion laws derived from the fundamental equations. By ascending from the three-dimensional physical space to the infinite-dimensional loop space, the inherent nonlinearity of the original advection problem is avoided, rendering it solvable with linear methods, albeit requiring less familiar mathematical techniques. Metaphorically, we mounted the dragon of loop space calculus and flew over the ’ice wall’ of nonlinearity.
Why discrete shells survive at fixed Sc
We now explain, in standard scaling terms, why discrete shells appear and persist at fixed $\mathit{Sc}$. In the decaying, high–Re limit we take $\nu\!\to\!0$ and $D\!\to\!0$ at fixed $\mathrm{Sc}=\nu/D$. Kelvin cancellation leaves a closed loop equation whose nonlinear balance sustains finite vorticity and a finite advective speed (of order unity), whereas the diffusive speed scales as $v_{D}\sim\sqrt{D/t}\ll 1$. Thus, advection dominates mixing, with diffusion acting as a secondary smoothener.
In this regime the scalar is transported ballistically to radii $r=\rho\,L(t)$, with $L(t)\sim\sqrt{t}$ and a dimensionless transport ratio $\rho$ that does not vanish as $\nu\!\to\!0$. The admissible values of $\rho$ form a discrete set fixed by a closure/periodicity constraint; for each denominator $q$, the number of admissible numerators (coprime to $q$) is $\varphi(q)$ (Euler’s totient). Because the advection–diffusion equation is linear in $T$, the ensemble superposes these discrete transports, producing concentric shells at $r=L(t)/n$ with parabolic profiles inside each shell and jumps proportional to $\varphi(n)/n^{2}$ across shells. Any finite $D$ rounds the discontinuities but leaves the discrete radii and the jump law intact at leading order. In Fourier variables, this structure appears as a universal function of $|\mathbf{q}|\,\big(\sqrt{t+t_{0}}-\sqrt{t_{0}}\big)$ with small, parameter–free oscillations.
A constructive realization of the admissible transport ratios is given in the momentum–loop formulation (see [Migdal(2025a)]), but it is not needed for the scaling argument above.
For a broader discussion relating these loop-space findings to classical turbulence paradigms, including Kolmogorov scaling, multifractal models, and the energy cascade narrative in the context of decaying turbulence, the reader is referred to Section 8 of our companion review [Migdal(2025a)].
The challenges of observing the spherical shells in DNS
The density distribution for a source, which starts as a delta function at the initial moment, is impossible to observe in DNS because the singularities cannot be resolved on the lattice. Even ignoring the delta function (which remains with some weaker weight at later stages of advection-diffusion), the lack of isotropy presents a statistical problem. Unlike correlation functions, which only depend on the difference between points in homogeneous isotropic turbulence, this density is not uniform, but rather tied to a specific point. To obtain statistics, one should run a parallel simulation in which the initial point is placed at random positions, and then average the results in relative coordinates. The practical alternative is to compute DNS statistics for the Fourier coefficients of the temperature. This transformation will provide effective averaging over the whole grid in coordinate space, adding statistics.
Besides, the predicted time dependence of the temperature amplitude in Fourier space is smooth, which would make it easier to compare with DNS data. There is a universal function depending on a scaling variable $\kappa=|\boldsymbol{q}|\tilde{\nu}(\sqrt{t+t_{0}}-\sqrt{t_{0}})$ . Therefore, by fitting $\tilde{\nu},t_{0}$ from the velocity spectrum, where the same parameters appear, one can pool all the statistics in the turbulent range of $|\boldsymbol{q}|$ for each time $t$ (the range found by studying the velocity spectrum). After this pooling, the resulting scaling function can be matched with our $U(\kappa)$ in (63).
Volume average of passive scalar
Figure 3: The plot of universal function $V(\xi)$ in (69) describing the average temperature as a function of the time $t$ as a function of scaling variable $\xi=\left(\sqrt{2\tilde{\nu}(t+t_{0})}-\sqrt{2\tilde{\nu}t_{0}}\right)/r_{0}$, where $r_{0}$ is the radius of the spherical averaging domain. Note the slight discontinuity of the derivative at $\xi=2,3,$..., when there is one more term added to the sum. This universal function is almost constant, because of the small coefficient $\frac{1}{240\pi^{2}}$.
Another way to improve statistics in DNS for passive scalars is to study volume averages. Let us consider the average temperature inside a spherical domain of radius $r_{0}$ centered at the initial point-like source. We have to take our point-like solution (54) and compute its mean value inside a spherical ball
$\bar{G}(r_{0},t)=\frac{3}{4\pi r_{0}^{3}}\int d^{3}\boldsymbol{r}\Theta(r_{0}-|\boldsymbol{r}|)G(r,t)=$
$\frac{3}{4\pi r_{0}^{3}}\left(1-\frac{\pi^{2}}{21600\zeta(5)}\right)+\frac{\pi^{2}}{2r_{0}^{3}}\int_{0}^{\frac{r_{0}}{2\pi L}}d\rho\rho^{4}\Phi\left(\left\lfloor\frac{1}{2\pi\rho}\right\rfloor\right);$ (64)
Using the definition of the Euler summatory function (49), we obtain the following series
$\bar{G}(r0,t)=\frac{3}{4\pi r_{0}^{3}}\left(1-\frac{\pi^{2}}{21600\zeta(5)}\right)+\frac{\pi^{2}}{2r_{0}^{3}}\sum_{k=1}^{\infty}\varphi(k)\int_{0}^{R}d\rho\rho^{4};$ (65)
$R=\min{\left(\frac{1}{2\pi k},\frac{r_{0}}{2\pi L}\right)}$ (66)
The integral terms can be arranged as follows
$\frac{\pi^{2}}{2r_{0}^{3}}\sum_{k=1}^{\infty}\varphi(k)\int_{0}^{R}d\rho\rho^{4}=\frac{\pi^{2}}{10r_{0}^{3}}\Theta(L-r_{0})\sum_{k=1}^{\lfloor\frac{L}{r_{0}}\rfloor}\varphi(k)\left(\left(\frac{r_{0}}{2\pi L}\right)^{5}-(2\pi k)^{-5}\right)+$
$\frac{\pi^{2}}{10r_{0}^{3}}\sum_{k=1}^{\infty}\varphi(k)(2\pi k)^{-5}$ (67)
The last sum reduces to the zeta function
$\frac{\pi}{28800r_{0}^{3}\zeta(5)}$ (68)
and cancels the similar term in $\bar{G}(r_{0},t)$. Collecting all terms, we find
$\bar{T}(r_{0},t)=\frac{T_{0}}{V_{0}}\left(1+\frac{t}{t_{0}}\right)^{-\frac{1}{2\mathit{Sc}}}V\left(\frac{\sqrt{2\tilde{\nu}(t+t_{0})}-\sqrt{2\tilde{\nu}t_{0}}}{r_{0}}\right);$ (69)
$V_{0}=\frac{4\pi r_{0}^{3}}{3};$ (70)
$V\left(\xi\right)=1+\frac{\Theta(\xi-1)}{240\pi^{2}}\sum_{k=1}^{\lfloor\xi\rfloor}\varphi(k)\left(\xi^{-5}-k^{-5}\right)$ (71)
This universal function $V(\xi)$ is still not mathematically smooth because of the floor value taken as an upper limit of the sum, but it is much smoother than the original shell structure. (see Figure 3). It is numerically very close to one, which would not be hard to verify in DNS.
Implications of the shell structure for the Millennium problem
The analytical solution for the passive scalar, detailed in Section 4(b) and illustrated in the figures above, carries profound implications for the broader Navier-Stokes Millennium problem. This solution, which describes the evolution of a scalar field coupled to the Navier-Stokes equations, presents an explicit, analytical solution with a finite-time singularity in the coupled NS-advection system.
This singularity is of a highly unusual kind. It is not a singular point, line of surface but rather a fractal set of concentric, spherical discontinuities (it is notable that Tao also conjectured a fractal singularity, albeit via a different mathematical mechanism [Tao(2016), Klarreich(2014)]). The scalar density profile is not smooth; it is composed of piecewise-parabolic profiles separated by finite, discontinuous jumps. In the limit, this intermittent structure condenses into a delta function.
Crucially, the geometric and statistical properties of this singular structure are dictated by number theory. The radial locations and jump magnitudes of the shells are governed by the Euler totient summatory function, which is rooted in the distribution of prime numbers. This finding provides the physical realization of a deep mathematical intuition, prophetically described by V.I. Arnol’d. Arnol’d identified that "The turbulence of number theory is the statistics of the values of the Euler function $\varphi(n)$… This is a very intermittent sequence…" [Arnol’d(1993)]. While Arnol’d proposed this as a mathematical analogy, the Euler ensemble solution provides the first-principles physical mechanism that derives this exact structure from the fundamental equations of fluid dynamics.
This framework thus unifies the concept of statistical intermittency with spatial singularities. The same underlying number-theoretic machinery that governs the spectrum of intermittency exponents (which we have linked to the zeros of the Riemann zeta function) also manifests spatially in the form of these quantized, discontinuous shells.
A critical open question remains regarding the stability of this solution. We have not proven that this singular state is a stable attractor for a generic set of initial conditions. However, its mere existence as an exact, analytical solution to the governing equations is significant. It demonstrates that the coupled NS-advection system permits non-perturbative, singular solutions. This finding lends support to conjectures, such as those by Tao, that singularities in nonlinear fluid equations are not only possible but may exhibit complex, fractal-like structures.
Disclaimer
We do not present a formal claim for the solution of the Millennium Problem. The arguments in this paper are presented in the language of theoretical physics, and no attempt is made to prove a rigorous mathematical theorem regarding this solution. Furthermore, the underlying Euler ensemble—while confirmed as an exact solution to the discretized loop equation [Elia and Lellis(2025)]—has not been rigorously proven to solve the Navier-Stokes equations in the continuum or under the specific smooth initial data conditions required by the Millennium Problem. From a formal mathematical standpoint, our results should be interpreted as conjectures. We hope they will inspire mathematicians to advance this theory to the level of rigor needed to establish a singular solution that satisfies the precise conditions of the Millennium Problem.
5 Conclusions
We have presented an exact analytic solution for the one–point scalar distribution from a point source in decaying turbulence, in the joint limit $\nu\to 0$, $D\to 0$ at fixed Schmidt number $\mathit{Sc}=\nu/D$. The background velocity statistics are those of the Euler ensemble. In this regime, the scalar is transported ballistically by the advective field while molecular smoothing is subleading; the mean field forms sharply defined, compactly supported spherical shells whose outer radius grows as $L(t)\sim\sqrt{t}$.
Main implications.
  • Closed scalar loop equation and analytic solution. The loop formulation yields a closed linear equation for the scalar loop functional and its analytic solution at fixed $\mathit{Sc}$. In real space, a point source produces a hierarchy of concentric shells at discrete radii $r=L(t)/n$, each with a piecewise–parabolic profile; the jump at the $n$th shell is proportional to $\varphi(n)/n^{2}$ (totient multiplicities). These structures form the underlying geometric structure of scalar transport in the decaying limit; any finite $D$ rounds the discontinuities but leaves the discrete radii at leading order.
  • Liquid vs static loop equivalence; dominance of advection. The Euler ensemble solves the liquid (Lagrangian) loop equation; on this solution, the advection term in the static (Eulerian) loop equation vanishes, so both ssequations hold. A scaling argument at fixed $\mathit{Sc}$ explains why the advective phase stays $O(1)$ while the diffusive factor is small, which is why the shells persist in the limit.
  • Fourier signature. For a monochromatic initial condition, the solution reduces to a universal, parameter–free response $U(\kappa)$ with an absolutely convergent series in $\kappa=|\mathbf{q}|\,(\sqrt{t+t_{0}}-\sqrt{t_{0}})$. The overall amplitude decays as $(1+t/t_{0})^{-1/(2\mathit{Sc})}$; $U(\kappa)$ exhibits small, predictable oscillations. This provides a readily measurable quantity in DNS that does not require resolving shells in real sspace.
  • Consistency with decaying–velocity statistics. The scalar results rest on the Euler ensemble, whose predictions for velocity decay and scales agree with classic experiments and modern DNS: turbulent kinetic energy $E(t)\sim t^{-5/4}$, $L(t)\sim t^{1/2}$, and a smoothly varying effective sindex for $\left\langle\Delta v^{2}\right\rangle(r)$ across the turbulent range (no plateau at $2/3$). No adjustable dimensionless parameters are used.
  • Verification strategy. Direct observation of shells may be challenging in finite-resolution flows, but the Fourier-space decay law and $U(\kappa)$ provide a robust test. In addition, the volume-averaged scalar density within a sphere yields a second, powerful real-space observable, $V(\xi)$, that is statistically smooth and directly verifiable in DNS.
  • Scope and regimes. The theory applies to unforced, decaying turbulence in the whole space, sin the high–Reynolds, vanishing–dissipation limit at fixed $\mathit{Sc}$. Weak forcing or finite diffusivity smooths the discontinuities; the discrete radii survive in the s preasymptotic regime. Uniqueness and stability of the attractor are interesting mathematical questions left open.
  • Early–time behaviour in DNS. For $t\ll t_{0}$, the factor $(1+t/t_{0})^{-1/(2\mathit{Sc})}$ is close to $\exp\left(-\frac{t}{2t_{0}\mathit{Sc}}\right)$, so long runs are beneficial for separating the algebraic decay from transient exponential behavior.
Outlook. Several extensions are natural: non–point initial data via the weight $W(\mathbf{q})$; two–point scalar statistics and fluxes; dependence on $\mathit{Sc}$ and weak inhomogeneity; stability of shells under weak forcing; and further rigor on the discrete–to–continuum limit (diffusion part) together with the momentum–space treatment of advection. On the validation side, targeted DNS that measures $U(\kappa)$ and, where feasible, the shell radii and jump law, would provide direct tests of the scalar predictions derived here. In particular, exploring the potential connection between these predicted shell structures and the experimentally observed ’ramp-cliff’ patterns in turbulent scalar fields [Sreenivasan(2019)] (see also Fig. 5 in [Migdal(2025a)]) warrants further investigation.
6 ACKNOWLEDGEMENTS
We gratefully acknowledge insightful discussions with Semon Rezchikov, Edoardo Spezzano, Katepalli Sreenivasan, and James Stone.
This research was partly supported by the Simons Foundation (award ID SFI-MPS-T-MPS-00010544) at the Institute for Advanced Study.
Declaration of Interests
The author reports no conflict of interest.
Declaration of AI use
I used AI-assisted writing tools (Gemini, GPT-5) solely for stylistic editing and text formatting of this manuscript. The tools were employed to improve readability and flow, standardize terminology and section headings, and convert punctuation to LaTeX-safe forms. No scientific content, derivations, equations, symbols, figures, data analysis, or results were generated or altered by the tool. I manually reviewed and verified all edited text for accuracy and consistency with the submitted manuscript. Any errors remain my responsibility.